### v1

# Betelgeuse, Betelgeuse, Betelgeuse, Betel-buddy? Constraints on the dynamical companion to  Orionis from HST

Abstract
Recently, two independent analyses have asserted that the cause of the Long Secondary Period (LSP) observed in the variability spectrum of our nearest red supergiant, Betelgeuse ( Ori), is an as-yet undetected, low-mass binary companion dubbed . In this paper, we present the results of a far-UV observational campaign using the STIS echelle spectrograph on the Hubble Space Telescope aimed at detecting spectral signatures of the companion. The four-quadrant tiling pattern and timing of the observations were optimized to isolate the companion, with observations taking place during a period of maximum angular and velocity separation between Betelgeuse and the putative companion. Spectral differencing between quadrants recovers no spectral features at the companionâs velocity in excess of the background or Betelgeuseâs chromosphere, i.e. a non-detection. Having determined that is most likely a Young Stellar Object (YSO) thanks to constraints from a complementary X-ray campaign with the Chandra X-ray Observatory in a companion paper, comparison of our data against canonical spectra from YSOs in the ULLYSES database allows us to confidently exclude masses above Mâ and companion continuum or line emission in excess of âergâs-1âcm-2âĂ -1 in the FUV (1200-1700Ă ). Future observational campaigns aware of the LSP phase are needed to place deeper constraints on the spectroscopic nature of .
1 Introduction
Alpha Orionis ( Ori), also known as Betelgeuse, is the 10th brightest star in the sky (Hoffleit & Jaschek, 1982; Hoffleit & Warren, 1987) and arguably our nearest red supergiant (RSG), with modern distance estimates ranging from 168âpc (Joyce et al., 2020) to 222âpc (Harper et al., 2017) from Earth. Interest in this star has increased significantly over the past five years due in large part to a severe brightness drop of âmag in 2019 that has come to be known as âThe Great Dimmingâ (e.g. Dupree et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Harper et al., 2020a, b; Kravchenko et al., 2021; Taniguchi et al., 2022; Cannon et al., 2023; JadlovskĂœ et al., 2024; Drevon et al., 2024). Given that Betelgeuse is an evolved massive star (, e.g. Joyce et al. 2020), this event prompted speculation that the star was close to undergoing a supernova (SN) explosion, but investigation suggests that Betelgeuseâs behavior during the Great Dimming comes from the creation of a dust cloud and a subsequent mode transition (Levesque & Massey, 2020; MontargĂšs et al., 2021; Dupree et al., 2022; MacLeod et al., 2023). These phenomena occur in the stellar and circumstellar envelope, and therefore need not be connected to any late-stage core activity that would suggest an imminent death. Current consensus favors an SN being several hundred thousand years away (e.g. Joyce et al., 2020; NeuhĂ€user et al., 2022; MacLeod et al., 2023), but tension in the literature nonetheless remains (Saio et al., 2023).
At the heart of the argument for or against an imminent SN is the classification of Betelgeuseâs periodicities and their driving mechanisms (Joyce et al., 2020; MacLeod et al., 2023). Betelgeuse exhibits a number of strong periodic signals, the most prominent of which are periodicities at roughly 2100 days, 400 days, and 185 days. If the 2100-day periodicity is an intrinsic variation corresponding to Betelgeuseâs fundamental (acoustic) pressure mode (FM), this would imply a radius in excess of and place the star in its core carbon fusion stage of evolution: a short-lived phase in the starâs last few hundred to few thousand years of life (Saio et al., 2023). The consensus regarding mode identification, however, is that the 400âday periodicity is the FM (Kiss et al., 2006; Joyce et al., 2020; Dupree et al., 2022; NeuhĂ€user et al., 2022; Granzer et al., 2022; MacLeod et al., 2023; MolnĂĄr et al., 2023) which represents the slowest (lowest-frequency) form of intrinsic variability caused by acoustic pressure variation.
This identification, coupled with other high-fidelity constraints, requires that Betelgeuse is undergoing He fusion in the core and therefore has âyr before its eventual SN. However, it also leaves open the question of what drives the longer (even lower-frequency), 2100âday periodicity. Long periodicities occurring at sub-FM frequencies are observed in about one third of cool, luminous stars with large, radially extended convective envelopes (e.g. Payne-Gaposchkin, 1954; Wood et al., 1999, 2004; Nicholls et al., 2009; SoszyĆski, 2007; SoszyĆski et al., 2009, 2011, 2013, 2021; Pawlak, 2021, 2023), including Red Supergiants such as Betelgeuse (e.g. Kiss et al., 2006; Yang & Jiang, 2012; Percy & Khatu, 2014; Chatys et al., 2019).
These sub-FM periodicities have been given the name âLong Secondary Periods,â or LSPs. There have been many mechanisms proposed to explain LSPs in general, ranging from giant cell convection (e.g. Stothers, 1972, 2010), mode interactions (e.g. Saio et al., 2015), non-radial pulsations (e.g. Wood, 2000), and dust modulation beyond the stellar surface (e.g. Wood et al., 2004; Wood & Nicholls, 2009; Percy & Shenoy, 2023). The most salient hypothesis, however, is binarity: namely, that the long secondary period is induced by an orbital companion (e.g. Wood et al. 2004; Derekas et al. 2006; SoszyĆski 2007; SoszyĆski et al. 2021).
Recently, Goldberg et al. (2024) and MacLeod et al. (2025) independently proposed the existence of a low-mass companion in a âd orbit as the explanation for the LSP observed in Betelgeuse, which is present in both the starâs lightcurve and its spectroscopic radial velocity (RV). Through a joint analysis of Betelgeuseâs brightness and RV variations and the phase offset between them compared to a sample of other stars with measured LSPs and RVs, Goldberg et al. (2024) argued that the binarity hypothesis is the best explanation for Betelgeuseâs LSP and RV signal, systematically ruling out all other proposed causes of the LSP in Betelgeuse. MacLeod et al. (2025) infer the existence of an using a long baseline of RV data as well as astrometric analysis, finding a periodic signal consistent with orbital motion on the LSP timescale in both data sets. In both of these works, the LSP is attributed to a low-mass () binary companion on a âyr, âAU, low-eccentricity orbit, at relatively low inclination relative to the observer. A companion at such a low mass and wide orbital separation would not directly affect Betelgeuseâs evolution through any substantial mass transfer, but it may interact indirectly via tides or by accreting from Betelgeuseâs stellar wind.
The idea that Betelgeuse could presently be a binary system has been explored in the literature. For example, from speckle-imaging measurements Karovska et al. (1986) proposed that Betelgeuse has two high-eccentricity companions, including one at a separation of 0.06â (near that of ) but a period of 2.1 years.111This particular configuration was not confirmed in follow-up work with higher-resolution instruments (e.g. Wilson et al., 1992; Kervella et al., 2009; MontargĂšs et al., 2016), nor recovered in the lightcurve or RV signal (Joyce et al., 2020; Goldberg et al., 2024; MacLeod et al., 2023, 2025). Moreover, a companion orbiting Betelgeuse is consistent with expectations for stars in Betelgeuseâs mass class (). Betelgeuse evolved from a late-O/early-B, main-sequence (MS) star. The binary fraction of late-O/early-B stars is 95% (e.g. Sana et al., 2012; de Mink et al., 2013). Though close binaries ( AU, which would have interacted by the RSG phase) show a uniform distribution of mass ratio with an excess at , at wider separations the mass ratio companions are skewed toward small mass ratios, favoring at 10AU (Rizzuto et al., 2013; Moe & Di Stefano, 2017; Offner et al., 2023). Other systems have been confirmed to have companions with low and intermediate separations, such as a large population of Cepheids in the range with binary companions on 1-20 yr periods, identified via both RV measurements (e.g. Evans et al., 2015) and X-ray emission and UV excess (Evans et al., 2022). It is therefore unsurprising that such a massive star has a companion at âAU.
Moving beyond dynamical arguments, observational follow-up informed by the companionâs orbital configuration will further constrain its nature and impact on the star and its environment. Importantly, Goldberg et al. (2024) show that the phase offset between the RV and light curves requires the companion to be behind Betelgeuse during Betelgeuseâs LSP brightness minima. The Great Dimming, which prompted many immediate follow-up observations, occurred at the brightness minimum of both the stellar pulsation and the LSP. However, at the time the companion was behind the disc of Betelgeuse, so observations then (e.g. Dupree et al., 2020, 2022; Kashyap et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Kravchenko et al., 2021; Alexeeva et al., 2021; Matthews & Dupree, 2022; Taniguchi et al., 2022) would not show any emission from the companion. Epochs corresponding to other orbital configurations, on the other hand, are ripe for further exploration.
MacLeod et al. (2025) further argue that the companion could plausibly induce rotational velocities of km/s via tidal interactions, matching the measured 36 year period in the lightcurve and RV (Joyce et al., 2020; MacLeod et al., 2025) as well as observations of a dipolar ârotationalâ velocity field in spatially-resolved observations (redshifted in the SE and blueshifted in the NW, see, e.g. Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018). Fast rotation in Betelgeuse has previously been argued to indicate a recent merger, as rapid rotation of a RSG at the km/s level requires an external source of angular momentum input into the starâs outer envelope (e.g. Wheeler et al., 2017; Wheeler & Chatzopoulos, 2023; Chatzopoulos et al., 2020; Nance et al., 2018; Shiber et al., 2024).222See, however, discussion by Ma et al. (2024) on under-resolved large-scale convection as a plausible origin for the Kervella et al. (2018) velocity field, a counter-hypothesis which predicts a dipole orientation which changes stochastically and decoheres on a year-long convective overturn timescale. If the companion is instead responsible, then when combined with an orbital phase inferred from Betelgeuseâs RV, the direction of rotation (PA of 48 W. of S.; Kervella et al. 2018) would constrain the companionâs projected orbital plane and thus on-sky location relative to Betelgeuse.
We obtained Directorâs Discretionary Time (DDT) during Hubble Space Telescope (HST) Cycle 32 to observe Betelgeuse and its putative companion with the Space Telescope Imaging Spectrograph (STIS) instrument in the Far-UV (FUV) in a configuration never before observed. Though MacLeod et al. (2025) argue that the chromospheric emission from Betelgeuse would prevent detection of a companion, the mass rangeâwhich extends potentially up to permitted by the high-quality STELLA RVs (Goldberg et al., 2024) and consistent with the astrometric analysis of MacLeod et al. (2025)âincludes a number of scenarios that could be detectable in both the far-UV and X-ray bands (see our companion publication in OâGrady et al. 2025).
This work is organized as follows: In Section 2, we discuss motivations for our choices to observe the Betelgeuse system in the FUV and at quadrature. Section 3 then describes our HST configuration, tiling pattern, and observational timing. Section 4 presents our spectroscopic observations, and discusses the constraints we can place on the FUV emission from . We summarize and conclude in Section 5.
2 Expectations for UV emission
Our observing strategy is informed by the various mass estimates and constraints for , and the most likely evolutionary possibilities for its companion. In this section we discuss the range of expectations for the companionâs spectral properties, and emphasize the importance of observing at âs peak expected offset in redshift from Betelgeuse.
2.1 Possible Companion Scenarios
A mass range of â0.5â2 Mâ enables a range of possible identities for the companion. One astrophysical object satisfying the mass condition is a low-mass, âstandardâ star. Under co-natal and coeval assumptions for the system (Dolan et al., 2016; Joyce et al., 2020; NeuhĂ€user et al., 2022), such a star must in fact be a young stellar object, or YSO, on the basis of Betelgeuseâs age of roughly 10 Myr (Tognelli et al., 2011; Baraffe et al., 2015; HaemmerlĂ© et al., 2019; Amard et al., 2019). It is possible that the companion is not co-natal, but rather has been captured due to dynamical interactions, with the system then ejected from its birth cluster. However, Betelgeuseâs proper motion has not yet been traced back to a specific cluster (though see some discussion in Harper et al., 2008; Bouy et al., 2014; Bouy & Alves, 2015). Moreover, ejections generally happen within 1 Myr of cluster formation (Oh & Kroupa, 2016), but the age of Betelgeuse and all nearby clusters and OB associations are much older than 1 Myr (e.g. Blaauw, 1964; Briceño et al., 2007, 2019), so even a dynamical ejection scenario would be close to coeval.
Other possibilities include a more exotic, compact object, such as a neutron star (NS) or white dwarf (WD). In the case of a WD, it is not possible to construct a non-interacting, two-body, coeval evolutionary scenario that combines a 10-Myr-old red supergiant with even the highest-mass WD, which has a theoretical lower age limit of 40 Myr and a practical (observed) lower age limit of 70 Myr. While more complex three-body (or more) scenarios involving the deposition of a WD into Betelgeuseâs orbit from some other system are theoretically possible, the kinematic signatures of such an event are unclear. Evolutionary timescales therefore preclude the most likely versions of a WD scenario.
Additionally, in the case of a compact object, we would expect to see a bright X-ray signature of accretion: given the orbital separation, the companion is embedded within the dusty shell surrounding Betelgeuse (e.g., Haubois et al., 2019, 2023), and would interact with Betelgeuseâs wind (Aizu, 1973; Livio et al., 1986; Blondin et al., 1990; Mukai, 2017; Maldonado et al., 2025). Typical X-ray luminosities for accreting WDs are erg s-1 (Luna et al., 2013), and for accreting NSs are erg s-1 (Yungelson et al., 2019; De et al., 2022a, b). Complementary observations with Chandra were performed within a few weeks of the HST observations and revealed no such signature, as discussed in detail in OâGrady et al. (2025).
An accreting compact object is therefore ruled out on the basis of a non-detection with Chandra over 40ks of observations: this integration time implies an upper limit of erg s-1 for the X-ray luminosity, orders of magnitude lower than the expected luminosity of erg s-1 for an accreting compact object (up to assumptions regarding the circumstellar hydrogen column depth). While a YSO would typically have an accretion disk as well, and therefore produce an X-ray signature, the wide diversity of X-ray luminosities from YSOs makes it impossible to place constraints on the stellar companion on the basis of the Chandra observations alone. It is also likely that, given the relatively small separation between Betelgeuse and and the fact that Betelgeuse was an OB star during its main sequence, any accretion disk on a YSO companion would have been irradiated away (see 2.3).
Therefore, if (even nearly) co-natal and coeval with Betelgeuse, within the mass parameters of either study is most likely a disk-free YSO, or pre-Main sequence (MS) star, of slightly later spectral class than its subsequent class at ZAMS.
2.2 Motivation to Observe at Quadrature
The luminosity ratio between Betelgeuse and in the YSO scenario is expected to be of order â, making undetectable in integrated light (MacLeod et al., 2025). However, in the far-UV spectrum of both YSOs and RSGs, emission from the chromosphere dominates the photosphere continuum (Linsky, 2017). HST has resolved many chromospheric lines in the spectra of YSOs (Yang et al., 2012; Roman-Duval et al., 2020, 2025), and has observed the chromosphere of Betelgeuse (Carpenter et al., 2018). This leads to three possible means of detection for : (1) the detection of a slight UV excess relative to observations taken where the companion is not present, (2) strong emission lines either shifted relative to, or not expected from, the chromosphere of Betelgeuse, and (3) RV variations in those UV lines over the timescale of an LSP period. Option (3) would require continuous observation over the 5.5-year duration of an LSP, which is impractical. As MacLeod et al. (2025) point out, option (1) is tenuous given intrinsic variations in Betelgeueseâs bright chromosphere. However, if observed at a proper orbital configuration where âs emission is redshifted or blueshifted relative to Betelgeuse, option (2) could be promising (whether or not the companion is coeval with Betelgeuse, and whether or not it is indeed a YSO).
We therefore target our observations such that âs emission is expected to be significantly redshifted relative to Betelgeuseâs, near receding quadrature on 6 December 2024 (Goldberg et al. 2024, within of the MacLeod et al. (2025) ephemeris which places quadrature slightly earlier in mid-to-late November 2024). We also note that the companionâs next (approaching) quadrature is years down the line (late November 2027) after passing behind the disc of Betelgeuse.
2.3 Far UV Expectations for a YSO Companion
In the far ultra-violet (FUV), the overall flux of Betelgeuse is expected to dominate over that of a YSO companion. Figure 1 shows the Color-Magnitude Diagram (CMD) for all Gaia stars with GALEX FUV photometry (â1,750âĂ ) across galactic latitudes â and longitudes â within âpc, which encompasses the 8â15âMyr old Scorpius-Centaurus (Sco-Cen) OB association. The official Sco-Cen members from Luhman (2022) are shown in blue. For both main sequence (MS) and 10âMyr pre-MS stars (excluding the white dwarf sequence to the bottom left), the Gaia (BPâRP) color roughly corresponds to stellar mass, indicated on the top x-axis. For (A/F stars), the field and Sco-Cen stars follow the same MS relation. G-type () and K-type () stars exhibit FUV excess compared to their field MS counterparts. Early M-stars () exhibit a wide range of FUV fluxes and chromospheric activity, some of them even brighter than the GK stars. This factor-of-ten spread in FUV brightness is expected, as emission lines and FUV flux depend on a starâs chromospheric activity, which varies with rotation and time and environmental properties in addition to mass and age.
The weighted average FUV flux of Betelgeuse (135â175ânm) is âergâs-1âcm-2âĂ -1 (Carpenter et al., 2018), which corresponds to an apparent brightness of âmag and absolute brightness of âmag in the GALEX passband given the 168âpc distance to Betelgeuse (Joyce et al., 2020). Horizontal lines indicating 100%, 10%, and 1% of Betelgeuseâs FUV brightness are shown on the CMD (Fig. 1). Based on the available E140M spectrum of Betelgeuse from Carpenter et al. (2018), a companion with 30% the FUV flux of Betelgeuse should have already been detected.
For possible masses of in the range, the companionâs luminosity in the FUV is expected to be of that of Betelgeuse, making it difficult to detect a shift in the overall flux level. However, the chromospheres of stars and YSOs in the Far UV are dominated by emission lines (see Linsky 2017 for a review). As discussed above, targeting our observations near the companionâs receding quadrature corresponds to Betelgeuseâs minimal RV (maximum Betelgeuse blueshift), but the companionâs maximal RV (maximum redshift). This orbital configuration entails that the emission features of should be redshifted with respect to the same features of Betelgeuse, by a speed of 45.5 km/s. This is well above the detectable limit for HST-STIS with a good resolution grating (10 km/s). Therefore, for emission lines that are sufficiently strong in both the chromospheres of Betelgeuse and YSOs, a second, redshifted emission line should be visible.
For want of first-principles chromospheric modeling, reasonable proxies for the possible FUV spectra of are provided by the ULLYSES project, which has compiled, in a publicly available database, the HST UV spectroscopy of many pre-MS/T Tauri stars (Roman-Duval et al., 2020, 2025). To illustrate this, we choose five YSOs from the ULLYSES dataset333https://ullyses.stsci.edu/ullyses-targets-ttauri.html across the range of relevant masses and ages â CVSO 104 (0.77Mâ, 11 Myr), CVSO 165 (0.91Mâ, 11 Myr), MP Mus (1.1Mâ, 15 Myr), HD1353344B (1.5Mâ, 9 Myr), and HD104237 (2.0Mâ, 5 Myr). Later in this paper, we also use two other YSOs for comparative purposes â PDS 70 (0.76Mâ, 11 Myr), and AK Sco (1.52Mâ, 11Myr). The masses are obtained from the ULLYSES website and the ages from estimates of the age of the cluster containing the YSO (Blaauw, 1964; de Geus et al., 1989; Mamajek et al., 2002; MĂŒller et al., 2011; Bossini et al., 2019; Esposito et al., 2020). These objects additionally represent a wide range of .
These data, along with the HST/STIS E140M spectrum of Betelgeuse from Carpenter et al. (2018), are shown in Figure 2. At the time of the Carpenter et al. (2018) observations (conducted in 2011), was transiting Betelgeuse, nearly at the middle of transit (see also § 3.4). Thus, while the overall flux of each emission line would contain contributions from both Betelgeuse and , their relative radial velocities should be close to zero â the only RV shift present at that time should be the overall peculiar motion of the system (22 km/s) as well as any RV changes induced by Betelgeuseâs pulsational or convective activity. In Figure 2, these data are shown as a thick solid orange line (labeled  Ori 2011), and have been blueshifted by km/s to match the expected total RV of Betelgeuse during our 2024 observations. We likewise shift the ULLYSES YSO spectra to be at the expected radial velocity of during the 2024 observations, 45.5 km/s redshifted. They have also been corrected for distance (168 pc, Joyce et al. 2020) and extinction ( and , MontargĂšs et al. 2021) appropriate to the parameters of the Betelgeuse system.
While the overall flux in the UV of these YSOs is only 1â5% of that of Betelgeuse, several features could be visible in a STIS exposure containing (as opposed to exposures containing only the chromosphere of Betelgeuse). There are 8 strong emission lines in the ULLYSES YSO spectra that are close to or as strong as the same line from Betelgeuse, and 4 that do not appear as emission lines in the spectrum of Betelgeuse. The four panels of Figure 2 highlight four wavelength ranges with promising features that may be visible in our observations of the system.444See Figure 7 and Table 2 of Linsky (2017) for the full suite of lines we examined here and in the observing proposal.
We also note that the ULLYSES data are all of isolated YSO systems. YSOs are often surrounded by, and accrete from, circumstellar disks, leading to the emission âwingsâ seen in some of the ULLYSES spectra. These are protoplanetary disks at early (10âMyr) times, and gas-poor processed disks at later (10âMyr) times (Ribas et al., 2015). These disks eventually photo-evaporate from high-energy radiation from the central star (Gorti & Hollenbach, 2009). However, these circumstellar disks can also be photoevaporated by UV radiation from nearby OB stars (Adams et al., 2004; Andrews, 2020), which has been directly observed for YSOs in young OB associations (Johnstone et al., 1998). For binary YSOs, systems with circumprimary but no circumsecondary disks are also much more common than the inverse (White & Ghez, 2001), and Kraus et al. (2012) finds that 2/3rds of YSO systems with close separations (r40âAU, compared to the 9AU separation for ) disperse their disks with 1âMyr of formation. Given that the  Oriâ system has a small separation relative to 40AU, and given that Betelguse was an extremely hot O star during its main-sequence lifetime, there is a high likelihood that a circumstellar disk around the companion has been dispersed by present day. Thus, the wide emission features stemming from strong winds or accretion (red/blue-shifted relative to the companion) that are common in YSO spectra around CII, and CIV lines (seen in the ULLYSES spectra in panels (a) and (d) of Figure 2), as well as wings around Lyman-, would almost certainly not be present in the  Ori system.
We therefore focus our efforts on narrow lines without those wings, such as fluorescent H2 lines (seen in all four panels of Figure 2), as well as Lyman (discussed in §4.3). Of the spectra shown here, the 2âMâ YSO would be plainly visible in these lines, while the 0.77âMâ YSO would never be visible even with a large RV offset between Betelgeuse and . The 0.91âMâ YSO also appears generally too faint to detect. The 1.1 and 1.5âMâ YSOs are more promising to identify, though the extra flux added to the system is small. We also note the important caveat that these are individual, empirical YSO observations, not theoretical models of YSO FUV chromospheric features across the full range of possible companion masses, for which first-principles calculations remain elusive.
3 Observations with HST
Below, we summarize the observational history of Betelgeuse to place our campaign in context. We then describe the parameters of our own observational campaign that took place in November 2024. We emphasize the importance of the timing of our observations relative to the companionâs orbital phase (receding quadrature). We also describe our unique aperture tiling pattern, which allows for co-incident spectra which both do and do not contain emission from the companion.
3.1 Past UV Observations
Long before the launch of HST, early measures of the ultraviolet spectrum of Betelgeuse (spectral type M1-M2) were obtained from a balloon-borne telescope-spectrometer, the OAOâ2 spacecraft (Wisconsin photometer-spectrometer), and OAOâ3 (Princeton experiment on Copernicus). All detected the Mg II doublet near 2800âĂ , a strong emission doublet in the near-ultraviolet (Kondo et al., 1972; Doherty, 1972; Bernat & Lambert, 1976). Subsequently, the International Ultraviolet Explorer (IUE) extended spectroscopy further into the ultraviolet to the Lyman- line near 1200âĂ . These low resolution spectra show no evidence of any species of ionization beyond Si II and Fe II, suggesting electron temperatures less than about 10,000âK (Basri et al., 1981). The Hopkins Ultraviolet Telescope on the Space Shuttle offered the possibility of spectroscopy from 800âĂ Â to 1800âĂ Â (Davidsen, 1993); however, the few spectra of Betelgeuse were too low signal-to-noise to be conclusive.
A substantial advance in the study of the near ultraviolet spectrum came with the Goddard High Resolution Spectrograph (GHRS) on HST, acquiring a detailed spectrum from 1980â3000âĂ Â that enabled the identification of many features (Brandt et al., 1995) from single and doubly ionized species, among them Fe II, C II, Ni II, Ti II in addition to molecular emission of OH. The GHRS instrument also accessed the far ultraviolet spectrum, from 1200âĂ Â to 1930âĂ , a region that is dominated by CO bandhead absorption in addition to the singly and doubly ionized species, C I, S I, O I, Cr II, and Si II (Carpenter et al., 1994). Those authors also claim the presence of a true continuum, with a color temperature of 3500â4800âK which they ascribe to the chromosphere, and suggest an origin in hot chromospheric regions. A large hot region was later observed on the extended ultraviolet disk in the 2500âĂ Â continuum, with temperatures in excess of 200âK from the disk value (Gilliland & Dupree, 1996), thus supporting origin in a hot region. Carpenter et al. (1994) in fact consider but dismiss the idea that a companion creates the continuum, in part because the observed flux distribution of warmer stars does not match the continuum shape.
The spectral range below 1200âĂ , became generally accessible with the launch of the Far Ultraviolet Spectroscopic Explorer (FUSE), which enabled access from 905â1179âĂ Â where highly ionized species are expected. However, a spectrum of Betelgeuse revealed no sign of C III or O VI which are found in giants and supergiants of spectral type F, G, and K (Dupree et al., 2005).
More recently, observations of Betelgeuse (PI T. Ayres 2010) were performed with STIS E140M using a square âarcsec aperture for 4.5 hours (Carpenter et al., 2018).The dates of these observations were February 7, 8, 24, 26, and April 1st 2011, at which point was transiting (passing in front of) Betelgeuse. While our campaign used the same instrument, grating, and aperture as the 2011 campaign, ours differs critically in timing, tiling, and motivation. The previous Ayres (2010); Carpenter et al. (2018) campaign was only interested in the overall UV spectrum of the chromosphere, and they did not vary their pointing. Therefore, that campaign did not obtain any spatial information and did not collect any âcleanâ spectra without .
3.2 Our FUV Observations and configuration-specific expectations
We obtained STIS E140M spectra at four different pointings with the 0.2â0.2â aperture, which we describe below in Section 3.3. Our campaign tiled the area including and immediately surrounding Betelgeuse into quadrants, as shown in Figure 3 with the expectation that (a point source, âmas) would be in one or at most two quadrants. A detection would therefore determine which quadrant Ori B is in, thereby constraining the geometry of the orbit, independent of inferences from Betelgeuseâs inferred rotation axis.
Calculations performed in the observing proposal demonstrated that, by subtracting two HST STIS E140M spectra, one with and one without , it would be possible to spot differences at the 2% level given our proposed exposure times and anticipated S/N. Likewise, the expected km/s RV shift given the systemâs orbital configuration is within the detectable limit for our chosen grating (10âkm/s for E140M); therefore, for strong emission lines from both stars, we also anticipated an apparent doubling of lines with a separation of 0.16Ă Â in the exposure that contains (see above discussion in §2.3). It is also worth noting that Betelgeuseâs chromospheric lines are substantially broadened by turbulence (km/s, Carpenter & Robinson 1997), which is expected to exceed the broadening of the low-mass companion, but would not exceed the companionâs orbital velocity. We predicted that our campaign would be sensitive to companions , similar to the expected mass of  recovered from the high-fidelity STELLA RV (Goldberg et al., 2024) and astrometric analysis (MacLeod et al., 2025). We also predicted sensitivity to young, chromospherically active 0.6 companions, consistent with the mass suggested by the detailed long-baseline RV analysis (MacLeod et al., 2025). Moreover, may be accreting at a higher rate due to focused winds from Betelgeuse and thus may be even brighter in the FUV than its isolated, 10âMyr-old counterparts in Sco-Cen. We therefore anticipated a reasonably good chance of detecting , and if not, placing constraints on remaining mass ranges allowed by a null detection.
3.3 HST configuration
Observations for this project were obtained using the STIS instrument on HST via a Cycle 32 Directorâs Discretionary award of 4 orbits (GO/DD-17881555stsci.edu/hst-program-info/program/?program=17881; PI: M. Joyce). The observations were separated into 4, 1-orbit visits in order to ease the scheduling burden introduced by the reduced gyro mode, and also to limit the number of repeat visits required should a failure occur. The first two visits executed on 2024 November 11 with the third and fourth visits occurring on 2024 November 15 and 25, respectively.
All four visits followed the same observing plan, which utilized a 1 second imaging target acquisition with the F25ND5 attenuating filter followed by 21000 second spectroscopic exposures using the E140M grating, 0.2â0.2â aperture, and a central wavelength of 1425 Ă . The central wavelength of 1425 Ă Â was chosen to ensure that key emission features from , such as Ly- (1215.67 Ă ), C IV (1548.2 Ă ), and fluorescent H2 lines (1446 Ă , 1505 Ă ) would be captured.
Since the low-mass companion is expected to be approximately 52âmas from the center of Betelgeuse, we adopted the tiling pattern shown in Fig. 3, which was constructed using combinations of Position Target (POS TARG) offset values of 0.084â in the X and Y coordinates. The four quadrants are labeled âNâ, âEâ, âSâ, and âWâ, loosely corresponding to their ordinal position on the sky. However, since no ORIENT constraints were applied, the roll angle of Visit 04 (the âWâ quadrant in Fig. 3) was slightly different than the previous three visits. While the exact quadrant in which is expected to appear is uncertain, the tiling strategy employed here incorporates sufficient overlap to guarantee its inclusion in at least one visit, if it is present.
If the orbital plane of the companion is indeed perpendicular to the PA rotation axis (Kervella et al., 2018) consistent with the rotation axis found by Uitenbroek et al. (1998); Lobel & Dupree (2001), the companion would appear in quadrant E just outside the overlap region with quadrant S, and in the SE stack. Likewise, insofar as the rotational velocity persists and extends into the chromosphere, emission from Betelgeuse itself should be redshifted in the E quadrant by a few km/s relative to the W quadrant.
3.4 Timing
Both the LSP phase and pulsation phase are important to interpret relative velocity and flux signatures in the UV spectra. As discussed above, the timing of our observations was targeted to match the receding quadrature of , near Betelgeuseâs RV minimum on the LSP timescale, approximately halfway between the maximum and minimum brightness phases of the LSP.
We compare the times of the HST observations to Betelgeuseâs light variations in Fig. 4. The two panels show 2000âd-long segments, i.e., covering about one LSP cycle with one intermediate segment between the two omitted from the plot. The 2011 observations analyzed by Carpenter et al. (2018), labeled Ayres+2010 for the observing proposal (Ayres, 2010), occurred near the LSPâs luminosity maximum on the LSP timescale, but while the pulsation was before and around minimum light. In contrast, our new observations (Joyce et al., 2024) were taken during the descending-luminosity phase of the LSP (broad behavior in the lower panel of Fig. 4, see also Fig. 5) and on the ascending-luminosity phase of the pulsation (local behavior in the lower panel of Fig. 4), just before maximum light.
Fig. 5 shows the timing of these same observations relative to the variations in Betelgeuseâs RV. The 2011 observations coincided with the STELLA RV observations, although the RV data are rather sparse at that time. We can subtract the average velocity and a sine fit approximating the 2170-day LSP variation from the STELLA RV data to estimate the contribution from stellar pulsations to the photospheric velocity. We find near-zero to low positive RV excess values that increase over time during the 2011 observations, indicating that the star was near maximum expansion, and then begun to shrink (increasing the redshift of the stellar photosphere, as can be seen in the upper panel of Fig. 5). Unfortunately, there are no simultaneous RV data publicly available for the new measurements, but we can infer the pulsation phase from the lightcurve. The starâs pulsation phase during our observations corresponds to the ascending branch of the light curve (seen in the lower panel of Fig. 4). The pulsation of the star is currently dominated by the O1 mode (MacLeod et al., 2023), for which Goldberg et al. (2024) found a phase lag between the photometry and the RV close to . This, and comparison with earlier simultaneous observations, suggest that the star itself was likely shrinking and approaching maximum compression at the time of our observations (Joyce et al., 2024). Therefore, we expect Betelgeuse to be somewhat cooler at the photosphere during the 2011 observations and hotter during our measurements, due to its underlying pulsational behavior.666Interestingly, the extremely high-velocity compression following the great dimming in April 2020 did not lead to significantly hotter , possibly because of the mass ejection (see, e.g. Wasatonic 2022 who show a long baseline of , TiO strength, and , as well as discussions by Levesque & Massey 2020). However, in general the pulsation phase does track the , seen also in the Wasatonic (2022) data. Likewise, Betelgeuseâs V-band magnitude is slightly fainter during the Joyce et al. (2024) observations as compared to Ayres (2010) (âmag as compared to âmag).
4 Analysis of Observations
We now discuss the similarities and differences between spectra taken in the four quadrants in the context of the search for the putative companion. For each of the four quadrants, the two exposures per visit were stacked using the stistools777https://stistools.readthedocs.io/en/latest/ and the HASP swrapper script (Debes et al., 2024). Flux values with SNR1 were then masked out in all figures showing spectra. For consistency, we apply the same stacking and masking procedure to the Ayres (2010) observations analyzed by Carpenter et al. (2018), and the same masking procedure to the ULLYSES data which we redshift to the companionâs expected location as in Figure 2.
4.1 Comparison of the 4 quadrants
Figure 6 shows our complete observations in all four quadrants with the same colors as in Figure 3, labeled by their approximate relative position on the sky. We also show the stacked north-west (NW) and south-east (SE) âhemispheresâ, also with the same colors as Figure 3. Each line is offset for legibility. The 2011 observations presented by Carpenter et al. (2018) are shown in black. Broadly, we see no systematic difference above the noise level between any of the four quadrants. Notable emission features have approximately the same width and height in all quadrants, and are comparable to the same features in the 2011 spectrum. The 2011 spectrum is slightly higher in flux at longer wavelengths, consistent with expectations that the stellar surface is slightly cooler due to the Betelgeuseâs underlying pulsational phase, and is less noisy than our observations due to the stacking of 6 overlapping exposures compared to our 2 exposures per quadrant (4 per hemisphere).
4.2 Emission from diagnostic lines
Figure 7 highlights specific wavelength ranges containing diagnostic lines useful for inferring the presence of a companion (see Figure 2 and surrounding discussion in §2.3). These four wavelength ranges represent the most promising areas (other than Ly-, see §4.3) to detect the companion. We plot the NW (cyan) and SE (magenta) stacked hemispheres of our observations888We show halves instead of quadrants for ease in viewing, but the results are the same when considering all four quadrants, the Carpenter et al. (2018) data (black), and several observations of YSOs from ULLYSES Roman-Duval et al. (2020, 2025) at ages and masses indicated in the legend (dark navy-to-green). These YSOs have been distance (168 pc Joyce et al. 2020)999At the farther astrometric distance of 222pc (Harper et al., 2017), this would lead to a factor of reduction in flux. and extinction ( and MontargÚs et al. 2021) corrected to the parameters of the Betelgeuse system.
First, we compare the shapes of emission features between the NW and SE hemispheres, and to the previous 2011 Carpenter et al. (2018) Betelgeuse observations. These three spectra represent three different system configurations: i) with emission only from Betelgeuse (likely the NW half), ii) with emission from the companion at a maximal RV shift of 45.5 km/s (likely the SE half), and iii) with emission from the companion at the same RV shift (transit) as Betelgeuse (Carpenter et al., 2018). We find no significant differences between the NW and SE halves. Further, with the Carpenter et al. (2018) data shifted to be at the approximate RV of our Betelgeuse observations, we find no differences between our data and that of Carpenter et al. (2018). All significant spectral features are the same to within noise limits. We thus report a clear non-detection of blueshifted emission features in excess of the chromospheric flux of Betelgeuse coming from a companion.
Given this, we look to the ULLYSES YSO data to illustrate what upper limits we can place on the mass and properties of the companion. We rule out the presence of a M2Mâ companion (light green line), which has flux greater than that of Betelgeuse in all cases. For the 0.77Mâ and 0.91Mâ YSOs (navy lines), even their chromospheric features with the highest flux relative to that of Betelgeuse (panel d) appear too low in flux to be discerned from noise in our spectra.
For the 1.1Mâ and 1.5Mâ YSOs, the presence of an RV-shifted feature in just one of our observed quadrants/halves would be difficult to disentangle from the noise (see, e.g. the CIV feature in panel d, as well as the CII doublet at 1334.5 and 1335.7 in panel a). In panels a), b), and c), for the narrow emission lines and not the disk emission features, we see that the flux of the 1.5Mâ YSO (e.g. the line near 1338.75 in panel a), or both that and the 1.1Mâ YSO (the CII lines near 1334.5-1336, H2 near 1505), meet or even slightly exceed that of all the Betelgeuse spectra.
If is a 1.1-1.5Mâ YSO, and has a spectral profile similar to the example YSOs from ULLYSES, then we would rule out those masses, though the diversity in emission from YSOs renders these ULLYSES spectra imperfect templates for the spectrum of the companion. Moreover, will be embedded within the dusty circumstellar material surrounding Betelgeuse, which may obscure or otherwise reprocess emission from the companion. Rather, these individual objects are illustrative at best, and their masses estimates.
While our observational campaign was designed to be maximally sensitive to a possible detection, given the diversity inherent in YSO spectra in the FUV, a non-detection is unfortunately only weakly constraining. Take, for example, the strong difference in overall flux in emission features between the 0.91Mâ and 1.1Mâ YSO spectra. Differences of up to an order of magnitude are visible (e.g. panel a around 1336.5), despite only 0.2Mâ between the mass estimates. Therefore, we cannot conclusively rule out the presence of a 1Mâ companion, though a smaller (M1Mâ) is favored.
4.3 Ly- Emission and Differences by quadrant
The Ly- (1215.67) emission line is another strong feature in the UV chromosphere of Betelgeuse as well as other stars and YSOs (Linsky, 2017; Carpenter et al., 2018; Roman-Duval et al., 2020, 2025). However, there is significant contamination of this feature from geocoronal emission from Earthâs atmosphere. The only HST UV observations of the Ly- line in a star uncontaminated by geocoronal emission are from stars with extremely high (150 km/s) radial velocities (Guinan et al., 2016; Schneider et al., 2019). The 22 km/s radial velocity of the Betelgeuse system is within the approximate width (35 km/s) of the observed Ly- line.
In an attempt to separate the source and geocoronal Ly- emission profiles, we extracted various cross-dispersion slices of the 2D spectra, including regions centered on the order traces where Ly- appears along with adjacent background regions on either side. Ly- appears in two adjacent orders in our E140M spectra, and for one of the orders the center trace spectrum shows evidence of a roughly 2.5 km s-1 shift relative to the background Ly- spectrum. However, no such signal was detected in the other order.
In the left panel of Figure 8, we plot the four quadrants of our Betelgeuse observations (same colors as Figure 3) along with the 2011 ASTRAL observations in black, shifted to the RV of Betelgeuse at the time of our 2024 observations. We see no evidence of an additional, RV-shifted Ly- line in any of the four quadrants. The Ly- emission has the same general shape and width (0.1) in all quadrants. We do observe a small shift in the center wavelengths of the emission feature â the east quadrant is offset from the west quadrant by km/s, while the north and south are offset from the west by approximately half that value. The implied orientation â with the eastern side of Betelgeuse redshifted relative to the western side â is consistent with the rotation axis inferred by Dupree et al. (1987); Gilliland & Dupree (1996); Uitenbroek et al. (1998); Lobel & Dupree (2001); Kervella et al. (2009, 2018), and others.
While large-scale convection ubiquitous in RSG envelopes (e.g. Stothers, 1972; Goldberg et al., 2022; Chiavassa et al., 2011, 2024; Ma et al., 2025) could over short timescales mimic both a changing astrometric signal (Chiavassa et al., 2020) and rotation inferred from an apparent redshift-blueshift dipole in resolved interferometric RV observations (Ma et al., 2024), convection is a stochastic process with decoherence times of order the envelope sound-crossing time (yr). The timing of our observations, on the other hand, has no correlation with the underlying convective phase or plume geometry of Betelgeuse. Therefore, the alignment of a south-westward RV dipole, persistent across many decades of observations and across many wavelengths including our 2024 observations, favors the claim that there is indeed coherent rotation of the near-surface layers.
When comparing our observations to the 2011 observations of Carpenter et al. (2018), we also find a surprising km/s offset between our median Ly- measurements and the 2011 Ly- measurement. This is the only emission feature throughout the entire spectrum that displays this offset between the 2024 and 2011 observations, including the 1302-1306 OI lines, which are also contaminated by geocoronal emission. All other features are only offset by 2.5 km/s, the RV difference in Betelgeuse between these dates, not offset by this km/s blueshift. We cannot explain this offset, but note that our 2024 data do show Ly- at approximately the same RV as other features in the spectrum.
Finally, in the right panel of Figure 8, we present our NW and SE spectra alongside several ULLYSES (Roman-Duval et al., 2020, 2025) YSOs, again distance and extinction corrected and shifted to âs RV at this phase. The YSOs plotted are different than in Figure 7; here we plot only ULLYSES YSOs observed with STIS, as COS is even more impacted by the geocoronal contamination at Ly- than STIS.
Though we were not able to extract the true shape of the Ly- emission unique to Betelgeuse, we do note that the total RV of relative to Earth is 65 km/s (the vertical orange dotted line in Figure 8), and is beyond the extent of the Ly- line. As seen in the right panel of Fig. 8, all reference YSO spectra (, , ) other than the 0.76Mâ YSO, the Ly- meet or exceed the flux of our Betelgeuse observations.
While the lack of a feature at the location of the Ly- line for the RV of is clear, we cannot firmly conclude that Ly- emission from a YSO with mass Mâ would be visible above the noise level of our 2024 spectra, as the Ly- features of the ULLYSES data are themselves dominated by geocoronal contamination. If we nonetheless expect comparable intrinsic Ly- emission from YSOs, then the lack of a second emission line here is constraining. However, without uncontaminated emission profiles, it is impossible to place an exact mass constraint using Ly-.
4.4 Implications for the Companion
Based on our comparison of observations of the Betelgeuse system during this campaign to characteristic spectral features from YSOs with mass determinations in the ULLYSES database, we confidently exclude a mass of for a YSO companion. We do not definitively exclude a mass of , but find the likelihood to be marginal. Masses below are not excluded. Figure 9 illustrates these scenarios with stellar evolutionary tracks (left panel) from MESA Isocrones and Stellar Tracks (MIST; Choi et al., 2016; Dotter, 2016; Paxton et al., 2011, 2013, 2015, 2018, left panel) at the masses given for the ULLYSES objects shown in Figure 7. The tracks shown span the Hayashi track through truncation at twice Betelgeuseâs approximate age, 20 Myr (prior to the zero-age main sequence in most cases). Lines of constant age (isochrones, red) at 5 Myr and 15 Myr are shown, with thick curves between (purple-blue-green) liberally indicating approximate plausible ages of . The region bounded by the red isochrones and purple-blue-green evolutionary tracks thus approximately illustrates the marginal (dashed) or fully consistent (solid) remaining locus of the companion (with redder and fainter companions than those shown also fully consistent).
The mass resolution of the ULLYSES observations we compare to, and more importantly, the extreme diversity of YSO luminosity in UV and high-energy emission (Preibisch et al. 2005; Preibisch & Feigelson 2005; Getman et al. 2005 and seen in our Figure 1), make it difficult to place strong constraints on the companion mass. Nevertheless, it is interesting that we only encounter true exclusion for a relatively high mass of (dotted green curve), and marginal constraints for the 1.1â1.5 cases (dashed blue-green curves). The fact that a YSO as massive as is not fully excluded leaves open the need for future resolved and pan-chromatic observing campaigns to place more detailed constraints on the companionâs physical nature and interaction with its environment, or, to identify a different cause of the Long Secondary Period persistent in Betelgeuseâs photometry, astrometry, and radial velocity from which the putative companion was inferred.
Likewise, flux-calibrated libraries like XSL (Chen et al., 2014; Manara et al., 2017) extended to the FUV would be valuable for constraining quiet or near-MS companions in the range which arenât well-captured in the ULLYSES data set, as broader libraries across a wider parameter space and absolute flux calibration could facilitate a more realistic detection threshold. Either way, the most promising observational path forward for direct detection remains long-baseline observations and analysis of historical observations of the resolved stellar surface and circumstellar envelope, specifically informed by timing relative to the LSP phase. Likewise, future and proposed FUV missions such as the Large UV/Optical/IR Surveyor (LUVOIR, The LUVOIR Team 2019) and the Cosmological Advanced Survey Telescope for Optical and ultraviolet Research (CASTOR, CĂŽte et al. 2012) may provide additional, deeper constraints, in particular on the UV continuum emission of Betelgeuse as well as higher-contrast features from the companion at favorable orbital configurations.
5 Conclusions
In this work, we detailed our HST STIS E140M observations of Betelgeuse during mid-to-late November 2024, near its recently-predicted (Goldberg et al., 2024; MacLeod et al., 2025) companionâs receding limb passage. This resulted in a non-detection: we do not find any statistically significant line emission in the FUV which could be identified as originating from the companion at the expected RV offset from Betelgeuse (km/s redshift, corresponding to a AU orbit with a d period), within flux limits of ergâs-1âcm-2âĂ -1 which increase with increasing wavelength (across the full band, the average flux is ergâs-1âcm-2âĂ -1, and the error is ergâs-1âcm-2âĂ -1). There is likewise no statistically significant contribution to continuum emission from a companion (i.e. the continuum differences between quadrants are consistent with HSTâs error), and continuum-like emission from Betelgeuse is just barely at the level compared to the error. Comparing to a set of YSO spectra taken with STIS in the ULLYSES database, we determine that higher-mass scenarios for the companion are ruled out, placing a spectroscopic mass constraint of and favoring masses , with acknowledgment of the great diversity in YSO spectra even within a given mass or spectral-type bin. This remains consistent with the lower-mass end of estimates fit from the STELLA RV only (Goldberg et al., 2024), but favors masses in the range as predicted by the long-baseline RV analysis (MacLeod et al., 2025).
Although we cannot uniquely identify spectroscopic features from the companion itself, we see slight differences in the spectra for each quadrant of our HST observations which loosely correspond on the sky to the Northern, Southern, Eastern, and Western quadrants of Betelgeuseâs chromosphere and circumstellar envelope. Though these differences are only slightly above the background, they likely indicate spatial variation across the chromosphere and circumstellar envelope. This is expected given the large-scale convective nature of RSG envelopes (see, e.g., recent review by Chiavassa et al., 2024), and is consistent with whatâs seen in interferometric studies of Betelgeuse (e.g. Harper et al., 2001; OâGorman et al., 2015; MontargĂšs et al., 2014, 2021; Kervella et al., 2009, 2018; Haubois et al., 2009, 2019, 2023; Kravchenko et al., 2021) as well as other work on spatially resolved UV observations of Betelgeuse and its chromosphere (e.g. Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Harper & Brown, 2006; Dupree et al., 2005, 2020; Matthews & Dupree, 2022). These differences are especially visible in the Ly- line (1215.67Ă ), though this feature is convolved with intense geocoronal emission and we were not able to cleanly separate the stellar Ly- profile. While the line itself is broadened by 10s of km/s, the velocity offsets of the Ly- line peaks in the four quadrants are consistent with the smeared-and-projected km/s rotational velocity (e.g. Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018). These spatial RV variations are also seen in the other emission features. Further modeling and future spatially-resolved observations may connect spectroscopic differences to spatial differences in chromospheric emission from Betelgeuse.
Additional X-ray observations with Chandra, presented in OâGrady et al. (2025), exclude a compact object (white dwarf or neutron star), and constrain the X-ray emission coming from a possible YSO companion with larger uncertainties arising from uncertainty in the X-ray absorption along the line-of-sight as well as the multiple-orders-of-magnitude differences in X-ray emission from YSOs even of the same spectral class. Therefore, future observational campaigns remain necessary for direct detection of , in order to confirm its presence and place constraints on its spectroscopic nature, interaction with the stellar wind, and any possible connection to dust formation such as during Great Dimming-like episodes. Finally, we re-emphasize the importance of considering the stellar variability cycles, and in particular Long Secondary Periods, when constructing spectroscopic campaigns for Betelgeuse and for all other binary and variable stars.
Data Availability
All HST data used in this work can be obtained from the MAST archive at http://dx.doi.org/10.17909/wpad-4s76 (catalog doi:10.17909/wpad-4s76).
Acknowledgements
We thank Morgan MacLeod for invaluable insight and discussions. We likewise acknowledge helpful conversations with Andrea Antoni, Matteo Cantiello, Thavisha Dharmawardena, Zarina Dhillon, Kareem El-Badry, Brian Metzger, Michael Palumbo, Adiv Paradise, Mathieu Renzo, and Lieke van Son. J.A.G. is supported by a Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. A.O. and B.O. gratefully acknowledge support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. M.J. gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commissionâs Widening Fellowship. This project has received funding from the European Unionâs Horizon 2020 research and innovation programme. This research was supported by the âSeismoLabâ KKP-137523 Ălvonal grant and by the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH). M.R.D. acknowledges support from the NSERC through grant RGPIN-2019-06186, the Ontario Ministry of Colleges and Universities through grant ER22-17-164, the Canada Research Chairs Program, and the Dunlap Institute at the University of Toronto.
We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). This research made use of NASAâs Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.

### v2

@@
 The four-quadrant tiling pattern and timing of the observations were optimized to isolate the companion, with observations taking place during a period of maximum angular and velocity separation between Betelgeuse and the putative companion.
-Spectral differencing between quadrants recovers no spectral features at the companionâs velocity in excess of the background or Betelgeuseâs chromosphere, i.e.
+Spectral differencing between quadrants recovers no spectral features at the companion’s velocity in excess of the background or Betelgeuse’s chromosphere, i.e.
 a non-detection.
-Having determined that is most likely a Young Stellar Object (YSO) thanks to constraints from a complementary X-ray campaign with the Chandra X-ray Observatory in a companion paper, comparison of our data against canonical spectra from YSOs in the ULLYSES database allows us to confidently exclude masses above Mâ and companion continuum or line emission in excess of âergâs-1âcm-2âĂ -1 in the FUV (1200-1700Ă ).
+Having determined that is most likely a Young Stellar Object (YSO) thanks to constraints from a complementary X-ray campaign with the Chandra X-ray Observatory in a companion paper, comparison of our data against canonical spectra from YSOs in the ULLYSES database allows us to confidently exclude masses above and companion continuum or line emission in excess of erg s-1 cm-2 Å-1 in the FUV (1200-1700Å).
 Future observational campaigns aware of the LSP phase are needed to place deeper constraints on the spectroscopic nature of .
 1 Introduction
-Alpha Orionis ( Ori), also known as Betelgeuse, is the 10th brightest star in the sky (Hoffleit & Jaschek, 1982; Hoffleit & Warren, 1987) and arguably our nearest red supergiant (RSG), with modern distance estimates ranging from 168âpc (Joyce et al., 2020) to 222âpc (Harper et al., 2017) from Earth.
-Interest in this star has increased significantly over the past five years due in large part to a severe brightness drop of âmag in 2019 that has come to be known as âThe Great Dimmingâ (e.g.
-Dupree et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Harper et al., 2020a, b; Kravchenko et al., 2021; Taniguchi et al., 2022; Cannon et al., 2023; JadlovskĂœ et al., 2024; Drevon et al., 2024).
+Alpha Orionis ( Ori), also known as Betelgeuse, is the 10th brightest star in the sky (Hoffleit & Jaschek, 1982; Hoffleit & Warren, 1987) and arguably our nearest red supergiant (RSG), with modern distance estimates ranging from 168 pc (Joyce et al., 2020) to 222 pc (Harper et al., 2017) from Earth.
+Interest in this star has increased significantly over the past five years due in large part to a severe brightness drop of mag in 2019 that has come to be known as “The Great Dimming” (e.g.
+Dupree et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; Montargès et al., 2021; Harper et al., 2020a, b; Kravchenko et al., 2021; Taniguchi et al., 2022; Cannon et al., 2023; Jadlovský et al., 2024; Drevon et al., 2024).
 Given that Betelgeuse is an evolved massive star (, e.g.
 Joyce et al.
-2020), this event prompted speculation that the star was close to undergoing a supernova (SN) explosion, but investigation suggests that Betelgeuseâs behavior during the Great Dimming comes from the creation of a dust cloud and a subsequent mode transition (Levesque & Massey, 2020; MontargĂšs et al., 2021; Dupree et al., 2022; MacLeod et al., 2023).
+2020), this event prompted speculation that the star was close to undergoing a supernova (SN) explosion, but investigation suggests that Betelgeuse’s behavior during the Great Dimming comes from the creation of a dust cloud and a subsequent mode transition (Levesque & Massey, 2020; Montargès et al., 2021; Dupree et al., 2022; MacLeod et al., 2023).
 These phenomena occur in the stellar and circumstellar envelope, and therefore need not be connected to any late-stage core activity that would suggest an imminent death.
 Current consensus favors an SN being several hundred thousand years away (e.g.
-Joyce et al., 2020; NeuhĂ€user et al., 2022; MacLeod et al., 2023), but tension in the literature nonetheless remains (Saio et al., 2023).
-At the heart of the argument for or against an imminent SN is the classification of Betelgeuseâs periodicities and their driving mechanisms (Joyce et al., 2020; MacLeod et al., 2023).
+Joyce et al., 2020; Neuhäuser et al., 2022; MacLeod et al., 2023), but tension in the literature nonetheless remains (Saio et al., 2023).
+At the heart of the argument for or against an imminent SN is the classification of Betelgeuse’s periodicities and their driving mechanisms (Joyce et al., 2020; MacLeod et al., 2023).
 Betelgeuse exhibits a number of strong periodic signals, the most prominent of which are periodicities at roughly 2100 days, 400 days, and 185 days.
-If the 2100-day periodicity is an intrinsic variation corresponding to Betelgeuseâs fundamental (acoustic) pressure mode (FM), this would imply a radius in excess of and place the star in its core carbon fusion stage of evolution: a short-lived phase in the starâs last few hundred to few thousand years of life (Saio et al., 2023).
-The consensus regarding mode identification, however, is that the 400âday periodicity is the FM (Kiss et al., 2006; Joyce et al., 2020; Dupree et al., 2022; NeuhĂ€user et al., 2022; Granzer et al., 2022; MacLeod et al., 2023; MolnĂĄr et al., 2023) which represents the slowest (lowest-frequency) form of intrinsic variability caused by acoustic pressure variation.
-This identification, coupled with other high-fidelity constraints, requires that Betelgeuse is undergoing He fusion in the core and therefore has âyr before its eventual SN.
-However, it also leaves open the question of what drives the longer (even lower-frequency), 2100âday periodicity.
+If the 2100-day periodicity is an intrinsic variation corresponding to Betelgeuse’s fundamental (acoustic) pressure mode (FM), this would imply a radius in excess of and place the star in its core carbon fusion stage of evolution: a short-lived phase in the star’s last few hundred to few thousand years of life (Saio et al., 2023).
+The consensus regarding mode identification, however, is that the 400–day periodicity is the FM (Kiss et al., 2006; Joyce et al., 2020; Dupree et al., 2022; Neuhäuser et al., 2022; Granzer et al., 2022; MacLeod et al., 2023; Molnár et al., 2023) which represents the slowest (lowest-frequency) form of intrinsic variability caused by acoustic pressure variation.
+This identification, coupled with other high-fidelity constraints, requires that Betelgeuse is undergoing He fusion in the core and therefore has yr before its eventual SN.
+However, it also leaves open the question of what drives the longer (even lower-frequency), 2100–day periodicity.
 Long periodicities occurring at sub-FM frequencies are observed in about one third of cool, luminous stars with large, radially extended convective envelopes (e.g.
-Payne-Gaposchkin, 1954; Wood et al., 1999, 2004; Nicholls et al., 2009; SoszyĆski, 2007; SoszyĆski et al., 2009, 2011, 2013, 2021; Pawlak, 2021, 2023), including Red Supergiants such as Betelgeuse (e.g.
+Payne-Gaposchkin, 1954; Wood et al., 1999, 2004; Nicholls et al., 2009; Soszyński, 2007; Soszyński et al., 2009, 2011, 2013, 2021; Pawlak, 2021, 2023), including Red Supergiants such as Betelgeuse (e.g.
 Kiss et al., 2006; Yang & Jiang, 2012; Percy & Khatu, 2014; Chatys et al., 2019).
-These sub-FM periodicities have been given the name âLong Secondary Periods,â or LSPs.
+These sub-FM periodicities have been given the name “Long Secondary Periods,” or LSPs.
 There have been many mechanisms proposed to explain LSPs in general, ranging from giant cell convection (e.g.
@@
 2004; Derekas et al.
-2006; SoszyĆski 2007; SoszyĆski et al.
+2006; Soszyński 2007; Soszyński et al.
 2021).
@@
 (2024) and MacLeod et al.
-(2025) independently proposed the existence of a low-mass companion in a âd orbit as the explanation for the LSP observed in Betelgeuse, which is present in both the starâs lightcurve and its spectroscopic radial velocity (RV).
-Through a joint analysis of Betelgeuseâs brightness and RV variations and the phase offset between them compared to a sample of other stars with measured LSPs and RVs, Goldberg et al.
-(2024) argued that the binarity hypothesis is the best explanation for Betelgeuseâs LSP and RV signal, systematically ruling out all other proposed causes of the LSP in Betelgeuse.
+(2025) independently proposed the existence of a low-mass companion in a d orbit as the explanation for the LSP observed in Betelgeuse, which is present in both the star’s lightcurve and its spectroscopic radial velocity (RV).
+Through a joint analysis of Betelgeuse’s brightness and RV variations and the phase offset between them compared to a sample of other stars with measured LSPs and RVs, Goldberg et al.
+(2024) argued that the binarity hypothesis is the best explanation for Betelgeuse’s LSP and RV signal, systematically ruling out all other proposed causes of the LSP in Betelgeuse.
 MacLeod et al.
 (2025) infer the existence of an using a long baseline of RV data as well as astrometric analysis, finding a periodic signal consistent with orbital motion on the LSP timescale in both data sets.
-In both of these works, the LSP is attributed to a low-mass () binary companion on a âyr, âAU, low-eccentricity orbit, at relatively low inclination relative to the observer.
-A companion at such a low mass and wide orbital separation would not directly affect Betelgeuseâs evolution through any substantial mass transfer, but it may interact indirectly via tides or by accreting from Betelgeuseâs stellar wind.
+In both of these works, the LSP is attributed to a low-mass () binary companion on a yr, AU, low-eccentricity orbit, at relatively low inclination relative to the observer.
+A companion at such a low mass and wide orbital separation would not directly affect Betelgeuse’s evolution through any substantial mass transfer, but it may interact indirectly via tides or by accreting from Betelgeuse’s stellar wind.
 The idea that Betelgeuse could presently be a binary system has been explored in the literature.
 For example, from speckle-imaging measurements Karovska et al.
-(1986) proposed that Betelgeuse has two high-eccentricity companions, including one at a separation of 0.06â (near that of ) but a period of 2.1 years.111This particular configuration was not confirmed in follow-up work with higher-resolution instruments (e.g.
-Wilson et al., 1992; Kervella et al., 2009; MontargĂšs et al., 2016), nor recovered in the lightcurve or RV signal (Joyce et al., 2020; Goldberg et al., 2024; MacLeod et al., 2023, 2025).
-Moreover, a companion orbiting Betelgeuse is consistent with expectations for stars in Betelgeuseâs mass class ().
+(1986) proposed that Betelgeuse has two high-eccentricity companions, including one at a separation of 0.06” (near that of ) but a period of 2.1 years.111This particular configuration was not confirmed in follow-up work with higher-resolution instruments (e.g.
+Wilson et al., 1992; Kervella et al., 2009; Montargès et al., 2016), nor recovered in the lightcurve or RV signal (Joyce et al., 2020; Goldberg et al., 2024; MacLeod et al., 2023, 2025).
+Moreover, a companion orbiting Betelgeuse is consistent with expectations for stars in Betelgeuse’s mass class ().
 Betelgeuse evolved from a late-O/early-B, main-sequence (MS) star.
@@
 Evans et al., 2015) and X-ray emission and UV excess (Evans et al., 2022).
-It is therefore unsurprising that such a massive star has a companion at âAU.
-Moving beyond dynamical arguments, observational follow-up informed by the companionâs orbital configuration will further constrain its nature and impact on the star and its environment.
+It is therefore unsurprising that such a massive star has a companion at AU.
+Moving beyond dynamical arguments, observational follow-up informed by the companion’s orbital configuration will further constrain its nature and impact on the star and its environment.
 Importantly, Goldberg et al.
-(2024) show that the phase offset between the RV and light curves requires the companion to be behind Betelgeuse during Betelgeuseâs LSP brightness minima.
+(2024) show that the phase offset between the RV and light curves requires the companion to be behind Betelgeuse during Betelgeuse’s LSP brightness minima.
 The Great Dimming, which prompted many immediate follow-up observations, occurred at the brightness minimum of both the stellar pulsation and the LSP.
 However, at the time the companion was behind the disc of Betelgeuse, so observations then (e.g.
-Dupree et al., 2020, 2022; Kashyap et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; MontargĂšs et al., 2021; Kravchenko et al., 2021; Alexeeva et al., 2021; Matthews & Dupree, 2022; Taniguchi et al., 2022) would not show any emission from the companion.
+Dupree et al., 2020, 2022; Kashyap et al., 2020; Dharmawardena et al., 2020; Levesque & Massey, 2020; Montargès et al., 2021; Kravchenko et al., 2021; Alexeeva et al., 2021; Matthews & Dupree, 2022; Taniguchi et al., 2022) would not show any emission from the companion.
 Epochs corresponding to other orbital configurations, on the other hand, are ripe for further exploration.
 MacLeod et al.
-(2025) further argue that the companion could plausibly induce rotational velocities of km/s via tidal interactions, matching the measured 36 year period in the lightcurve and RV (Joyce et al., 2020; MacLeod et al., 2025) as well as observations of a dipolar ârotationalâ velocity field in spatially-resolved observations (redshifted in the SE and blueshifted in the NW, see, e.g.
+(2025) further argue that the companion could plausibly induce rotational velocities of km/s via tidal interactions, matching the measured 36 year period in the lightcurve and RV (Joyce et al., 2020; MacLeod et al., 2025) as well as observations of a dipolar “rotational” velocity field in spatially-resolved observations (redshifted in the SE and blueshifted in the NW, see, e.g.
 Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018).
-Fast rotation in Betelgeuse has previously been argued to indicate a recent merger, as rapid rotation of a RSG at the km/s level requires an external source of angular momentum input into the starâs outer envelope (e.g.
+Fast rotation in Betelgeuse has previously been argued to indicate a recent merger, as rapid rotation of a RSG at the km/s level requires an external source of angular momentum input into the star’s outer envelope (e.g.
 Wheeler et al., 2017; Wheeler & Chatzopoulos, 2023; Chatzopoulos et al., 2020; Nance et al., 2018; Shiber et al., 2024).222See, however, discussion by Ma et al.
@@
 (2018) velocity field, a counter-hypothesis which predicts a dipole orientation which changes stochastically and decoheres on a year-long convective overturn timescale.
-If the companion is instead responsible, then when combined with an orbital phase inferred from Betelgeuseâs RV, the direction of rotation (PA of 48 W.
+If the companion is instead responsible, then when combined with an orbital phase inferred from Betelgeuse’s RV, the direction of rotation (PA of 48 W.
 of S.; Kervella et al.
-2018) would constrain the companionâs projected orbital plane and thus on-sky location relative to Betelgeuse.
-We obtained Directorâs Discretionary Time (DDT) during Hubble Space Telescope (HST) Cycle 32 to observe Betelgeuse and its putative companion with the Space Telescope Imaging Spectrograph (STIS) instrument in the Far-UV (FUV) in a configuration never before observed.
+2018) would constrain the companion’s projected orbital plane and thus on-sky location relative to Betelgeuse.
+We obtained Director’s Discretionary Time (DDT) during Hubble Space Telescope (HST) Cycle 32 to observe Betelgeuse and its putative companion with the Space Telescope Imaging Spectrograph (STIS) instrument in the Far-UV (FUV) in a configuration never before observed.
 Though MacLeod et al.
-(2025) argue that the chromospheric emission from Betelgeuse would prevent detection of a companion, the mass rangeâwhich extends potentially up to permitted by the high-quality STELLA RVs (Goldberg et al., 2024) and consistent with the astrometric analysis of MacLeod et al.
-(2025)âincludes a number of scenarios that could be detectable in both the far-UV and X-ray bands (see our companion publication in OâGrady et al.
+(2025) argue that the chromospheric emission from Betelgeuse would prevent detection of a companion, the mass range—which extends potentially up to permitted by the high-quality STELLA RVs (Goldberg et al., 2024) and consistent with the astrometric analysis of MacLeod et al.
+(2025)—includes a number of scenarios that could be detectable in both the far-UV and X-ray bands (see our companion publication in O’Grady et al.
 2025).
@@
 Our observing strategy is informed by the various mass estimates and constraints for , and the most likely evolutionary possibilities for its companion.
-In this section we discuss the range of expectations for the companionâs spectral properties, and emphasize the importance of observing at âs peak expected offset in redshift from Betelgeuse.
+In this section we discuss the range of expectations for the companion’s spectral properties, and emphasize the importance of observing at ’s peak expected offset in redshift from Betelgeuse.
 2.1 Possible Companion Scenarios
-A mass range of â0.5â2 Mâ enables a range of possible identities for the companion.
-One astrophysical object satisfying the mass condition is a low-mass, âstandardâ star.
-Under co-natal and coeval assumptions for the system (Dolan et al., 2016; Joyce et al., 2020; NeuhĂ€user et al., 2022), such a star must in fact be a young stellar object, or YSO, on the basis of Betelgeuseâs age of roughly 10 Myr (Tognelli et al., 2011; Baraffe et al., 2015; HaemmerlĂ© et al., 2019; Amard et al., 2019).
+A mass range of 0.5–2 M⊙ enables a range of possible identities for the companion.
+One astrophysical object satisfying the mass condition is a low-mass, “standard” star.
+Under co-natal and coeval assumptions for the system (Dolan et al., 2016; Joyce et al., 2020; Neuhäuser et al., 2022), such a star must in fact be a young stellar object, or YSO, on the basis of Betelgeuse’s age of roughly 10 Myr (Tognelli et al., 2011; Baraffe et al., 2015; Haemmerlé et al., 2019; Amard et al., 2019).
 It is possible that the companion is not co-natal, but rather has been captured due to dynamical interactions, with the system then ejected from its birth cluster.
-However, Betelgeuseâs proper motion has not yet been traced back to a specific cluster (though see some discussion in Harper et al., 2008; Bouy et al., 2014; Bouy & Alves, 2015).
+However, Betelgeuse’s proper motion has not yet been traced back to a specific cluster (though see some discussion in Harper et al., 2008; Bouy et al., 2014; Bouy & Alves, 2015).
 Moreover, ejections generally happen within 1 Myr of cluster formation (Oh & Kroupa, 2016), but the age of Betelgeuse and all nearby clusters and OB associations are much older than 1 Myr (e.g.
@@
 In the case of a WD, it is not possible to construct a non-interacting, two-body, coeval evolutionary scenario that combines a 10-Myr-old red supergiant with even the highest-mass WD, which has a theoretical lower age limit of 40 Myr and a practical (observed) lower age limit of 70 Myr.
-While more complex three-body (or more) scenarios involving the deposition of a WD into Betelgeuseâs orbit from some other system are theoretically possible, the kinematic signatures of such an event are unclear.
+While more complex three-body (or more) scenarios involving the deposition of a WD into Betelgeuse’s orbit from some other system are theoretically possible, the kinematic signatures of such an event are unclear.
 Evolutionary timescales therefore preclude the most likely versions of a WD scenario.
-Additionally, in the case of a compact object, we would expect to see a bright X-ray signature of accretion: given the orbital separation, the companion is embedded within the dusty shell surrounding Betelgeuse (e.g., Haubois et al., 2019, 2023), and would interact with Betelgeuseâs wind (Aizu, 1973; Livio et al., 1986; Blondin et al., 1990; Mukai, 2017; Maldonado et al., 2025).
+Additionally, in the case of a compact object, we would expect to see a bright X-ray signature of accretion: given the orbital separation, the companion is embedded within the dusty shell surrounding Betelgeuse (e.g., Haubois et al., 2019, 2023), and would interact with Betelgeuse’s wind (Aizu, 1973; Livio et al., 1986; Blondin et al., 1990; Mukai, 2017; Maldonado et al., 2025).
 Typical X-ray luminosities for accreting WDs are erg s-1 (Luna et al., 2013), and for accreting NSs are erg s-1 (Yungelson et al., 2019; De et al., 2022a, b).
-Complementary observations with Chandra were performed within a few weeks of the HST observations and revealed no such signature, as discussed in detail in OâGrady et al.
+Complementary observations with Chandra were performed within a few weeks of the HST observations and revealed no such signature, as discussed in detail in O’Grady et al.
 (2025).
@@
 2.2 Motivation to Observe at Quadrature
-The luminosity ratio between Betelgeuse and in the YSO scenario is expected to be of order â, making undetectable in integrated light (MacLeod et al., 2025).
+The luminosity ratio between Betelgeuse and in the YSO scenario is expected to be of order –, making undetectable in integrated light (MacLeod et al., 2025).
 However, in the far-UV spectrum of both YSOs and RSGs, emission from the chromosphere dominates the photosphere continuum (Linsky, 2017).
@@
 As MacLeod et al.
-(2025) point out, option (1) is tenuous given intrinsic variations in Betelgeueseâs bright chromosphere.
-However, if observed at a proper orbital configuration where âs emission is redshifted or blueshifted relative to Betelgeuse, option (2) could be promising (whether or not the companion is coeval with Betelgeuse, and whether or not it is indeed a YSO).
-We therefore target our observations such that âs emission is expected to be significantly redshifted relative to Betelgeuseâs, near receding quadrature on 6 December 2024 (Goldberg et al.
+(2025) point out, option (1) is tenuous given intrinsic variations in Betelgeuese’s bright chromosphere.
+However, if observed at a proper orbital configuration where ’s emission is redshifted or blueshifted relative to Betelgeuse, option (2) could be promising (whether or not the companion is coeval with Betelgeuse, and whether or not it is indeed a YSO).
+We therefore target our observations such that ’s emission is expected to be significantly redshifted relative to Betelgeuse’s, near receding quadrature on 6 December 2024 (Goldberg et al.
 2024, within of the MacLeod et al.
 (2025) ephemeris which places quadrature slightly earlier in mid-to-late November 2024).
-We also note that the companionâs next (approaching) quadrature is years down the line (late November 2027) after passing behind the disc of Betelgeuse.
+We also note that the companion’s next (approaching) quadrature is years down the line (late November 2027) after passing behind the disc of Betelgeuse.
 2.3 Far UV Expectations for a YSO Companion
 In the far ultra-violet (FUV), the overall flux of Betelgeuse is expected to dominate over that of a YSO companion.
-Figure 1 shows the Color-Magnitude Diagram (CMD) for all Gaia stars with GALEX FUV photometry (â1,750âĂ ) across galactic latitudes â and longitudes â within âpc, which encompasses the 8â15âMyr old Scorpius-Centaurus (Sco-Cen) OB association.
+Figure 1 shows the Color-Magnitude Diagram (CMD) for all Gaia stars with GALEX FUV photometry (–1,750 Å) across galactic latitudes – and longitudes – within pc, which encompasses the 8–15 Myr old Scorpius-Centaurus (Sco-Cen) OB association.
 The official Sco-Cen members from Luhman (2022) are shown in blue.
-For both main sequence (MS) and 10âMyr pre-MS stars (excluding the white dwarf sequence to the bottom left), the Gaia (BPâRP) color roughly corresponds to stellar mass, indicated on the top x-axis.
+For both main sequence (MS) and 10 Myr pre-MS stars (excluding the white dwarf sequence to the bottom left), the Gaia (BP–RP) color roughly corresponds to stellar mass, indicated on the top x-axis.
 For (A/F stars), the field and Sco-Cen stars follow the same MS relation.
@@
 Early M-stars () exhibit a wide range of FUV fluxes and chromospheric activity, some of them even brighter than the GK stars.
-This factor-of-ten spread in FUV brightness is expected, as emission lines and FUV flux depend on a starâs chromospheric activity, which varies with rotation and time and environmental properties in addition to mass and age.
-The weighted average FUV flux of Betelgeuse (135â175ânm) is âergâs-1âcm-2âĂ -1 (Carpenter et al., 2018), which corresponds to an apparent brightness of âmag and absolute brightness of âmag in the GALEX passband given the 168âpc distance to Betelgeuse (Joyce et al., 2020).
-Horizontal lines indicating 100%, 10%, and 1% of Betelgeuseâs FUV brightness are shown on the CMD (Fig.
-1).
+This factor-of-ten spread in FUV brightness is expected, as emission lines and FUV flux depend on a star’s chromospheric activity, which varies with rotation and time and environmental properties in addition to mass and age.
+The weighted average FUV flux of Betelgeuse (135–175 nm) is erg s-1 cm-2 Å-1 (Carpenter et al., 2018), which corresponds to an apparent brightness of mag and absolute brightness of mag in the GALEX passband given the 168 pc distance to Betelgeuse (Joyce et al., 2020).
+Horizontal lines indicating 100%, 10%, and 1% of Betelgeuse’s FUV brightness are shown on the CMD (Figure 1).
 Based on the available E140M spectrum of Betelgeuse from Carpenter et al.
 (2018), a companion with 30% the FUV flux of Betelgeuse should have already been detected.
-For possible masses of in the range, the companionâs luminosity in the FUV is expected to be of that of Betelgeuse, making it difficult to detect a shift in the overall flux level.
+For possible masses of in the range, the companion’s luminosity in the FUV is expected to be of that of Betelgeuse, making it difficult to detect a shift in the overall flux level.
 However, the chromospheres of stars and YSOs in the Far UV are dominated by emission lines (see Linsky 2017 for a review).
-As discussed above, targeting our observations near the companionâs receding quadrature corresponds to Betelgeuseâs minimal RV (maximum Betelgeuse blueshift), but the companionâs maximal RV (maximum redshift).
+As discussed above, targeting our observations near the companion’s receding quadrature corresponds to Betelgeuse’s minimal RV (maximum Betelgeuse blueshift), but the companion’s maximal RV (maximum redshift).
 This orbital configuration entails that the emission features of should be redshifted with respect to the same features of Betelgeuse, by a speed of 45.5 km/s.
@@
 For want of first-principles chromospheric modeling, reasonable proxies for the possible FUV spectra of are provided by the ULLYSES project, which has compiled, in a publicly available database, the HST UV spectroscopy of many pre-MS/T Tauri stars (Roman-Duval et al., 2020, 2025).
-To illustrate this, we choose five YSOs from the ULLYSES dataset333https://ullyses.stsci.edu/ullyses-targets-ttauri.html across the range of relevant masses and ages â CVSO 104 (0.77Mâ, 11 Myr), CVSO 165 (0.91Mâ, 11 Myr), MP Mus (1.1Mâ, 15 Myr), HD1353344B (1.5Mâ, 9 Myr), and HD104237 (2.0Mâ, 5 Myr).
-Later in this paper, we also use two other YSOs for comparative purposes â PDS 70 (0.76Mâ, 11 Myr), and AK Sco (1.52Mâ, 11Myr).
-The masses are obtained from the ULLYSES website and the ages from estimates of the age of the cluster containing the YSO (Blaauw, 1964; de Geus et al., 1989; Mamajek et al., 2002; MĂŒller et al., 2011; Bossini et al., 2019; Esposito et al., 2020).
+To illustrate this, we choose five YSOs from the ULLYSES dataset333https://ullyses.stsci.edu/ullyses-targets-ttauri.html across the range of relevant masses and ages – CVSO 104 (0.77M⊙, 11 Myr), CVSO 165 (0.91M⊙, 11 Myr), MP Mus (1.1M⊙, 15 Myr), HD1353344B (1.5M⊙, 9 Myr), and HD104237 (2.0M⊙, 5 Myr).
+Later in this paper, we also use two other YSOs for comparative purposes – PDS 70 (0.76M⊙, 11 Myr), and AK Sco (1.52M⊙, 11Myr).
+The masses, distances, and values are obtained from the ULLYSES website and the ages from estimates of the age of the cluster containing the YSO (Blaauw, 1964; de Geus et al., 1989; Mamajek et al., 2002; Müller et al., 2011; Bossini et al., 2019; Esposito et al., 2020).
 These objects additionally represent a wide range of .
@@
 (2018) observations (conducted in 2011), was transiting Betelgeuse, nearly at the middle of transit (see also § 3.4).
-Thus, while the overall flux of each emission line would contain contributions from both Betelgeuse and , their relative radial velocities should be close to zero â the only RV shift present at that time should be the overall peculiar motion of the system (22 km/s) as well as any RV changes induced by Betelgeuseâs pulsational or convective activity.
+Thus, while the overall flux of each emission line would contain contributions from both Betelgeuse and , their relative radial velocities should be close to zero — the only RV shift present at that time should be the overall peculiar motion of the system (22 km/s) as well as any RV changes induced by Betelgeuse’s pulsational or convective activity.
 In Figure 2, these data are shown as a thick solid orange line (labeled Ori 2011), and have been blueshifted by km/s to match the expected total RV of Betelgeuse during our 2024 observations.
 We likewise shift the ULLYSES YSO spectra to be at the expected radial velocity of during the 2024 observations, 45.5 km/s redshifted.
-They have also been corrected for distance (168 pc, Joyce et al.
-2020) and extinction ( and , MontargĂšs et al.
-2021) appropriate to the parameters of the Betelgeuse system.
-While the overall flux in the UV of these YSOs is only 1â5% of that of Betelgeuse, several features could be visible in a STIS exposure containing (as opposed to exposures containing only the chromosphere of Betelgeuse).
+We also correct for the distance of Betelgeuse (168 pc, Joyce et al.
+2020).
+To account for extinction, we de-extinct the YSO spectra given their values from the ULLYSES project, and re-extinct them with full extinction curves given the Betelgeuse’s measured extinction.
+Following Montargès et al.
+(2021) we adopt values of , appropriate for supergiants (Massey et al., 2005), and 444The of Betelgeuse has been largely consistent across observational works.
+Early work by Carpenter et al.
+(1994) fit the GHRS-FUV spectrum with =0.75 and =3.
+Levesque & Massey (2020) find =0.62 in their 2004-2020 observations.
+.
+This is likely a conservative estimate, as any denser gas and dust interior to the orbit of , such as the inner shell at inferred from ALMA observations (e.g.
+Perrin et al., 2007; Haubois et al., 2019, 2023), would contribute to the extinction of Betelgeuse but not .
+Extinction parameters are applied to the YSO spectra for the full Milky Way extinction curve (Gordon et al., 2023) using the dust_extinction555https://dust-extinction.readthedocs.io/en/stable/ package (Gordon, 2024).
+While the overall flux in the UV of these YSOs is only 1–5% of that of Betelgeuse, several features could be visible in a STIS exposure containing (as opposed to exposures containing only the chromosphere of Betelgeuse).
 There are 8 strong emission lines in the ULLYSES YSO spectra that are close to or as strong as the same line from Betelgeuse, and 4 that do not appear as emission lines in the spectrum of Betelgeuse.
-The four panels of Figure 2 highlight four wavelength ranges with promising features that may be visible in our observations of the system.444See Figure 7 and Table 2 of Linsky (2017) for the full suite of lines we examined here and in the observing proposal.
+The four panels of Figure 2 highlight four wavelength ranges with promising features that may be visible in our observations of the system.666See Figure 7 and Table 2 of Linsky (2017) for the full suite of lines we examined here and in the observing proposal.
 We also note that the ULLYSES data are all of isolated YSO systems.
-YSOs are often surrounded by, and accrete from, circumstellar disks, leading to the emission âwingsâ seen in some of the ULLYSES spectra.
-These are protoplanetary disks at early (10âMyr) times, and gas-poor processed disks at later (10âMyr) times (Ribas et al., 2015).
+YSOs are often surrounded by, and accrete from, circumstellar disks, leading to the emission “wings” seen in some of the ULLYSES spectra.
+These are protoplanetary disks at early (10 Myr) times, and gas-poor processed disks at later (10 Myr) times (Ribas et al., 2015).
 These disks eventually photo-evaporate from high-energy radiation from the central star (Gorti & Hollenbach, 2009).
@@
 For binary YSOs, systems with circumprimary but no circumsecondary disks are also much more common than the inverse (White & Ghez, 2001), and Kraus et al.
-(2012) finds that 2/3rds of YSO systems with close separations (r40âAU, compared to the 9AU separation for ) disperse their disks with 1âMyr of formation.
-Given that the  Oriâ system has a small separation relative to 40AU, and given that Betelguse was an extremely hot O star during its main-sequence lifetime, there is a high likelihood that a circumstellar disk around the companion has been dispersed by present day.
-Thus, the wide emission features stemming from strong winds or accretion (red/blue-shifted relative to the companion) that are common in YSO spectra around CII, and CIV lines (seen in the ULLYSES spectra in panels (a) and (d) of Figure 2), as well as wings around Lyman-, would almost certainly not be present in the  Ori system.
-We therefore focus our efforts on narrow lines without those wings, such as fluorescent H2 lines (seen in all four panels of Figure 2), as well as Lyman (discussed in §4.3).
-Of the spectra shown here, the 2âMâ YSO would be plainly visible in these lines, while the 0.77âMâ YSO would never be visible even with a large RV offset between Betelgeuse and .
-The 0.91âMâ YSO also appears generally too faint to detect.
-The 1.1 and 1.5âMâ YSOs are more promising to identify, though the extra flux added to the system is small.
+(2012) finds that 2/3rds of YSO systems with close separations (r40 AU, compared to the 9AU separation for ) disperse their disks with 1 Myr of formation.
+Given that the Ori– system has a small separation relative to 40AU, and given that Betelguse was an extremely hot O star during its main-sequence lifetime, there is a high likelihood that a circumstellar disk around the companion has been dispersed by present day.
+Thus, the wide emission features stemming from strong winds or accretion (red/blue-shifted relative to the companion) that are common in YSO spectra (seen in the ULLYSES spectra in panels (a) and (d) of Figure 2), as well as wings around Lyman-, would almost certainly not be present in the Ori system.
+We therefore focus our efforts on narrow lines without those wings, such as fluorescent H2 lines (seen in all four panels of Figure 2).
+Of the spectra shown here, the 2 M⊙ YSO would be plainly visible in these lines, while the 0.77 M⊙ YSO would never be visible even with a large RV offset between Betelgeuse and .
+The 0.91 M⊙ YSO also appears generally too faint to detect.
+The 1.1 and 1.5 M⊙ YSOs are more promising to identify, though the extra flux added to the system is small.
 We also note the important caveat that these are individual, empirical YSO observations, not theoretical models of YSO FUV chromospheric features across the full range of possible companion masses, for which first-principles calculations remain elusive.
@@
 We then describe the parameters of our own observational campaign that took place in November 2024.
-We emphasize the importance of the timing of our observations relative to the companionâs orbital phase (receding quadrature).
+We emphasize the importance of the timing of our observations relative to the companion’s orbital phase (receding quadrature).
 We also describe our unique aperture tiling pattern, which allows for co-incident spectra which both do and do not contain emission from the companion.
 3.1 Past UV Observations
-Long before the launch of HST, early measures of the ultraviolet spectrum of Betelgeuse (spectral type M1-M2) were obtained from a balloon-borne telescope-spectrometer, the OAOâ2 spacecraft (Wisconsin photometer-spectrometer), and OAOâ3 (Princeton experiment on Copernicus).
-All detected the Mg II doublet near 2800âĂ , a strong emission doublet in the near-ultraviolet (Kondo et al., 1972; Doherty, 1972; Bernat & Lambert, 1976).
-Subsequently, the International Ultraviolet Explorer (IUE) extended spectroscopy further into the ultraviolet to the Lyman- line near 1200âĂ .
-These low resolution spectra show no evidence of any species of ionization beyond Si II and Fe II, suggesting electron temperatures less than about 10,000âK (Basri et al., 1981).
-The Hopkins Ultraviolet Telescope on the Space Shuttle offered the possibility of spectroscopy from 800âĂ Â to 1800âĂ Â (Davidsen, 1993); however, the few spectra of Betelgeuse were too low signal-to-noise to be conclusive.
-A substantial advance in the study of the near ultraviolet spectrum came with the Goddard High Resolution Spectrograph (GHRS) on HST, acquiring a detailed spectrum from 1980â3000âĂ Â that enabled the identification of many features (Brandt et al., 1995) from single and doubly ionized species, among them Fe II, C II, Ni II, Ti II in addition to molecular emission of OH.
-The GHRS instrument also accessed the far ultraviolet spectrum, from 1200âĂ Â to 1930âĂ , a region that is dominated by CO bandhead absorption in addition to the singly and doubly ionized species, C I, S I, O I, Cr II, and Si II (Carpenter et al., 1994).
-Those authors also claim the presence of a true continuum, with a color temperature of 3500â4800âK which they ascribe to the chromosphere, and suggest an origin in hot chromospheric regions.
-A large hot region was later observed on the extended ultraviolet disk in the 2500âĂ Â continuum, with temperatures in excess of 200âK from the disk value (Gilliland & Dupree, 1996), thus supporting origin in a hot region.
+Long before the launch of HST, early measures of the ultraviolet spectrum of Betelgeuse (spectral type M1-M2) were obtained from a balloon-borne telescope-spectrometer, the OAO–2 spacecraft (Wisconsin photometer-spectrometer), and OAO–3 (Princeton experiment on Copernicus).
+All detected the Mg II doublet near 2800 Å, a strong emission doublet in the near-ultraviolet (Kondo et al., 1972; Doherty, 1972; Bernat & Lambert, 1976).
+Subsequently, the International Ultraviolet Explorer (IUE) extended spectroscopy further into the ultraviolet to the Lyman- line near 1200 Å.
+These low resolution spectra show no evidence of any species of ionization beyond Si II and Fe II, suggesting electron temperatures less than about 10,000 K (Basri et al., 1981).
+The Hopkins Ultraviolet Telescope on the Space Shuttle offered the possibility of spectroscopy from 800 Å to 1800 Å (Davidsen, 1993); however, the few spectra of Betelgeuse were too low signal-to-noise to be conclusive.
+A substantial advance in the study of the near ultraviolet spectrum came with the Goddard High Resolution Spectrograph (GHRS) on HST, acquiring a detailed spectrum from 1980–3000 Å that enabled the identification of many features (Brandt et al., 1995) from single and doubly ionized species, among them Fe II, C II, Ni II, Ti II in addition to molecular emission of OH.
+The GHRS instrument also accessed the far ultraviolet spectrum, from 1200 Å to 1930 Å, a region that is dominated by CO bandhead absorption in addition to the singly and doubly ionized species, C I, S I, O I, Cr II, and Si II (Carpenter et al., 1994).
+Those authors also claim the presence of a true continuum, with a color temperature of 3500–4800 K which they ascribe to the chromosphere, and suggest an origin in hot chromospheric regions.
+A large hot region was later observed on the extended ultraviolet disk in the 2500 Å continuum, with temperatures in excess of 200 K from the disk value (Gilliland & Dupree, 1996), thus supporting origin in a hot region.
 Carpenter et al.
 (1994) in fact consider but dismiss the idea that a companion creates the continuum, in part because the observed flux distribution of warmer stars does not match the continuum shape.
-The spectral range below 1200âĂ , became generally accessible with the launch of the Far Ultraviolet Spectroscopic Explorer (FUSE), which enabled access from 905â1179âĂ Â where highly ionized species are expected.
+The spectral range below 1200 Å, became generally accessible with the launch of the Far Ultraviolet Spectroscopic Explorer (FUSE), which enabled access from 905–1179 Å where highly ionized species are expected.
 However, a spectrum of Betelgeuse revealed no sign of C III or O VI which are found in giants and supergiants of spectral type F, G, and K (Dupree et al., 2005).
 More recently, observations of Betelgeuse (PI T.
-Ayres 2010) were performed with STIS E140M using a square âarcsec aperture for 4.5 hours (Carpenter et al., 2018).The dates of these observations were February 7, 8, 24, 26, and April 1st 2011, at which point was transiting (passing in front of) Betelgeuse.
+Ayres 2010) were performed with STIS E140M using a square arcsec aperture for 4.5 hours (Carpenter et al., 2018).The dates of these observations were February 7, 8, 24, 26, and April 1st 2011, at which point was transiting (passing in front of) Betelgeuse.
 While our campaign used the same instrument, grating, and aperture as the 2011 campaign, ours differs critically in timing, tiling, and motivation.
@@
 (2018) campaign was only interested in the overall UV spectrum of the chromosphere, and they did not vary their pointing.
-Therefore, that campaign did not obtain any spatial information and did not collect any âcleanâ spectra without .
+Therefore, that campaign did not obtain any spatial information and did not collect any “clean” spectra without .
 3.2 Our FUV Observations and configuration-specific expectations
-We obtained STIS E140M spectra at four different pointings with the 0.2â0.2â aperture, which we describe below in Section 3.3.
-Our campaign tiled the area including and immediately surrounding Betelgeuse into quadrants, as shown in Figure 3 with the expectation that (a point source, âmas) would be in one or at most two quadrants.
-A detection would therefore determine which quadrant Ori B is in, thereby constraining the geometry of the orbit, independent of inferences from Betelgeuseâs inferred rotation axis.
+We obtained STIS E140M spectra at four different pointings with the 0.2”0.2” aperture, which we describe below in Section 3.3.
+Our campaign tiled the area including and immediately surrounding Betelgeuse into quadrants, as shown in Figure 3 with the expectation that (a point source, mas) would be in one or at most two quadrants.
+A detection would therefore localize Ori B within a quadrant or hemisphere, thereby providing a constraint on the geometry of the orbit independent of inferences from Betelgeuse’s apparent rotation axis.
+In addition to the location of Betelgeuse and expected location of the companion if it is aligned with the Kervella et al.
+(2018) rotation axis, we indicate the diameter Point Spread Function (PSF) of STIS E140M as a horizontal bar.
+If the companion is present within the aperture of a given observation, then its observed image will be blurred over this range, with the majority of the flux remaining centrally concentrated.
+Flux from Betelgeuse itself, present in all 4 quadrants, will also be smeared.
+This provides additional motivation for quadrant-to-quadrant and hemisphere-to-hemisphere comparisons, rather than a finer-resolution search.
 Calculations performed in the observing proposal demonstrated that, by subtracting two HST STIS E140M spectra, one with and one without , it would be possible to spot differences at the 2% level given our proposed exposure times and anticipated S/N.
-Likewise, the expected km/s RV shift given the systemâs orbital configuration is within the detectable limit for our chosen grating (10âkm/s for E140M); therefore, for strong emission lines from both stars, we also anticipated an apparent doubling of lines with a separation of 0.16Ă Â in the exposure that contains (see above discussion in §2.3).
-It is also worth noting that Betelgeuseâs chromospheric lines are substantially broadened by turbulence (km/s, Carpenter & Robinson 1997), which is expected to exceed the broadening of the low-mass companion, but would not exceed the companionâs orbital velocity.
+Likewise, the expected km/s RV shift given the system’s orbital configuration is within the detectable limit for our chosen grating (10 km/s for E140M); therefore, for strong emission lines from both stars, we also anticipated an apparent doubling of lines with a separation of 0.16Å in the exposure that contains (see above discussion in §2.3).
+It is also worth noting that Betelgeuse’s chromospheric lines are substantially broadened by turbulence (km/s, Carpenter & Robinson 1997), which is expected to exceed the broadening of the low-mass companion, but would not exceed the companion’s orbital velocity.
 We predicted that our campaign would be sensitive to companions , similar to the expected mass of recovered from the high-fidelity STELLA RV (Goldberg et al., 2024) and astrometric analysis (MacLeod et al., 2025).
 We also predicted sensitivity to young, chromospherically active 0.6 companions, consistent with the mass suggested by the detailed long-baseline RV analysis (MacLeod et al., 2025).
-Moreover, may be accreting at a higher rate due to focused winds from Betelgeuse and thus may be even brighter in the FUV than its isolated, 10âMyr-old counterparts in Sco-Cen.
+Moreover, may be accreting at a higher rate due to focused winds from Betelgeuse and thus may be even brighter in the FUV than its isolated, 10 Myr-old counterparts in Sco-Cen.
 We therefore anticipated a reasonably good chance of detecting , and if not, placing constraints on remaining mass ranges allowed by a null detection.
 3.3 HST configuration
-Observations for this project were obtained using the STIS instrument on HST via a Cycle 32 Directorâs Discretionary award of 4 orbits (GO/DD-17881555stsci.edu/hst-program-info/program/?program=17881; PI: M.
+Observations for this project were obtained using the STIS instrument on HST via a Cycle 32 Director’s Discretionary award of 4 orbits (GO/DD-17881777stsci.edu/hst-program-info/program/?program=17881; PI: M.
 Joyce).
@@
 The first two visits executed on 2024 November 11 with the third and fourth visits occurring on 2024 November 15 and 25, respectively.
-All four visits followed the same observing plan, which utilized a 1 second imaging target acquisition with the F25ND5 attenuating filter followed by 21000 second spectroscopic exposures using the E140M grating, 0.2â0.2â aperture, and a central wavelength of 1425 Ă .
-The central wavelength of 1425 Ă Â was chosen to ensure that key emission features from , such as Ly- (1215.67 Ă ), C IV (1548.2 Ă ), and fluorescent H2 lines (1446 Ă , 1505 Ă ) would be captured.
-Since the low-mass companion is expected to be approximately 52âmas from the center of Betelgeuse, we adopted the tiling pattern shown in Fig.
-3, which was constructed using combinations of Position Target (POS TARG) offset values of 0.084â in the X and Y coordinates.
-The four quadrants are labeled âNâ, âEâ, âSâ, and âWâ, loosely corresponding to their ordinal position on the sky.
-However, since no ORIENT constraints were applied, the roll angle of Visit 04 (the âWâ quadrant in Fig.
-3) was slightly different than the previous three visits.
+All four visits followed the same observing plan, which utilized a 1 second imaging target acquisition with the F25ND5 attenuating filter followed by 21000 second spectroscopic exposures using using the broad-band E140M-1425 grating setting (1150-1700 Å) and 0.2”0.2” aperture.
+This choice ensured that key emission features from , such as Ly- (1215.67 Å), C IV (1548.2 Å), and fluorescent H2 lines (1446 Å, 1505 Å) would be captured.
+Since the low-mass companion is expected to be approximately 52 mas from the center of Betelgeuse, we adopted the tiling pattern shown in Figure 3, which was constructed using combinations of Position Target (POS TARG) offset values of 0.084” in the X and Y coordinates.
+The four quadrants are labeled “N”, “E”, “S”, and “W”, loosely corresponding to their ordinal position on the sky.
+However, since no ORIENT constraints were applied, the roll angle of Visit 04 (the “W” quadrant in Figure 3) was slightly different than the previous three visits.
 While the exact quadrant in which is expected to appear is uncertain, the tiling strategy employed here incorporates sufficient overlap to guarantee its inclusion in at least one visit, if it is present.
@@
 (1998); Lobel & Dupree (2001), the companion would appear in quadrant E just outside the overlap region with quadrant S, and in the SE stack.
-Likewise, insofar as the rotational velocity persists and extends into the chromosphere, emission from Betelgeuse itself should be redshifted in the E quadrant by a few km/s relative to the W quadrant.
 3.4 Timing
 Both the LSP phase and pulsation phase are important to interpret relative velocity and flux signatures in the UV spectra.
-As discussed above, the timing of our observations was targeted to match the receding quadrature of , near Betelgeuseâs RV minimum on the LSP timescale, approximately halfway between the maximum and minimum brightness phases of the LSP.
-We compare the times of the HST observations to Betelgeuseâs light variations in Fig.
-4.
-The two panels show 2000âd-long segments, i.e., covering about one LSP cycle with one intermediate segment between the two omitted from the plot.
+As discussed above, the timing of our observations was targeted to match the receding quadrature of , near Betelgeuse’s RV minimum on the LSP timescale, approximately halfway between the maximum and minimum brightness phases of the LSP.
+We compare the times of the HST observations to Betelgeuse’s light variations in Figure 4.
+The two panels show 2000 d-long segments, i.e., covering about one LSP cycle with one intermediate segment between the two omitted from the plot.
 The 2011 observations analyzed by Carpenter et al.
-(2018), labeled Ayres+2010 for the observing proposal (Ayres, 2010), occurred near the LSPâs luminosity maximum on the LSP timescale, but while the pulsation was before and around minimum light.
-In contrast, our new observations (Joyce et al., 2024) were taken during the descending-luminosity phase of the LSP (broad behavior in the lower panel of Fig.
-4, see also Fig.
-5) and on the ascending-luminosity phase of the pulsation (local behavior in the lower panel of Fig.
-4), just before maximum light.
-Fig.
-5 shows the timing of these same observations relative to the variations in Betelgeuseâs RV.
+(2018), labeled Ayres+2010 for the observing proposal (Ayres, 2010), occurred near the LSP’s luminosity maximum on the LSP timescale, but while the pulsation was before and around minimum light.
+In contrast, our new observations (Joyce et al., 2024) were taken during the descending-luminosity phase of the LSP (broad behavior in the lower panel of Figure 4, see also Figure 5) and on the ascending-luminosity phase of the pulsation (local behavior in the lower panel of Figure 4), just before maximum light.
+Figure 5 shows the timing of these same observations relative to the variations in Betelgeuse’s RV.
 The 2011 observations coincided with the STELLA RV observations, although the RV data are rather sparse at that time.
 We can subtract the average velocity and a sine fit approximating the 2170-day LSP variation from the STELLA RV data to estimate the contribution from stellar pulsations to the photospheric velocity.
-We find near-zero to low positive RV excess values that increase over time during the 2011 observations, indicating that the star was near maximum expansion, and then begun to shrink (increasing the redshift of the stellar photosphere, as can be seen in the upper panel of Fig.
-5).
+We find near-zero to low positive RV excess values that increase over time during the 2011 observations, indicating that the star was near maximum expansion, and then begun to shrink (increasing the redshift of the stellar photosphere, as can be seen in the upper panel of Figure 5).
 Unfortunately, there are no simultaneous RV data publicly available for the new measurements, but we can infer the pulsation phase from the lightcurve.
-The starâs pulsation phase during our observations corresponds to the ascending branch of the light curve (seen in the lower panel of Fig.
-4).
+The star’s pulsation phase during our observations corresponds to the ascending branch of the light curve (seen in the lower panel of Figure 4).
 The pulsation of the star is currently dominated by the O1 mode (MacLeod et al., 2023), for which Goldberg et al.
@@
 This, and comparison with earlier simultaneous observations, suggest that the star itself was likely shrinking and approaching maximum compression at the time of our observations (Joyce et al., 2024).
-Therefore, we expect Betelgeuse to be somewhat cooler at the photosphere during the 2011 observations and hotter during our measurements, due to its underlying pulsational behavior.666Interestingly, the extremely high-velocity compression following the great dimming in April 2020 did not lead to significantly hotter , possibly because of the mass ejection (see, e.g.
+Therefore, we expect Betelgeuse to be somewhat cooler at the photosphere during the 2011 observations and hotter during our measurements, due to its underlying pulsational behavior.888Interestingly, the extremely high-velocity compression following the great dimming in April 2020 did not lead to significantly hotter , possibly because of the mass ejection (see, e.g.
 Wasatonic 2022 who show a long baseline of , TiO strength, and , as well as discussions by Levesque & Massey 2020).
 However, in general the pulsation phase does track the , seen also in the Wasatonic (2022) data.
-Likewise, Betelgeuseâs V-band magnitude is slightly fainter during the Joyce et al.
-(2024) observations as compared to Ayres (2010) (âmag as compared to âmag).
+Likewise, Betelgeuse’s V-band magnitude is slightly fainter during the Joyce et al.
+(2024) observations as compared to Ayres (2010) ( mag as compared to mag).
 4 Analysis of Observations
 We now discuss the similarities and differences between spectra taken in the four quadrants in the context of the search for the putative companion.
-For each of the four quadrants, the two exposures per visit were stacked using the stistools777https://stistools.readthedocs.io/en/latest/ and the HASP swrapper script (Debes et al., 2024).
+For each of the four quadrants, the two exposures per visit were stacked using the stistools999https://stistools.readthedocs.io/en/latest/ and the HASP swrapper script (Debes et al., 2024).
 Flux values with SNR1 were then masked out in all figures showing spectra.
 For consistency, we apply the same stacking and masking procedure to the Ayres (2010) observations analyzed by Carpenter et al.
-(2018), and the same masking procedure to the ULLYSES data which we redshift to the companionâs expected location as in Figure 2.
+(2018), and the same masking procedure to the ULLYSES data which we redshift to the companion’s expected location as in Figure 2.
 4.1 Comparison of the 4 quadrants
 Figure 6 shows our complete observations in all four quadrants with the same colors as in Figure 3, labeled by their approximate relative position on the sky.
-We also show the stacked north-west (NW) and south-east (SE) âhemispheresâ, also with the same colors as Figure 3.
+We also show the stacked north-west (NW) and south-east (SE) ‘hemispheres’, also with the same colors as Figure 3.
 Each line is offset for legibility.
@@
 Notable emission features have approximately the same width and height in all quadrants, and are comparable to the same features in the 2011 spectrum.
-The 2011 spectrum is slightly higher in flux at longer wavelengths, consistent with expectations that the stellar surface is slightly cooler due to the Betelgeuseâs underlying pulsational phase, and is less noisy than our observations due to the stacking of 6 overlapping exposures compared to our 2 exposures per quadrant (4 per hemisphere).
+The 2011 spectrum is slightly higher in flux at longer wavelengths, consistent with expectations that the stellar surface is slightly cooler due to the Betelgeuse’s underlying pulsational phase, and is less noisy than our observations due to the stacking of 6 overlapping exposures compared to our 2 exposures per quadrant (4 per hemisphere).
+Individual spectral features in the different quadrants do exhibit some differences from each other.
+Figure 7 shows the region surrounding the H2 1504.82Å line and measurement uncertainties for the four quadrants as well as the two stacked hemispheres.
+Differences in the line shape coming from Betelgeuse itself are to be expected, and have been observed in other spatially resolved spectroscopic observations with HST (e.g.
+Uitenbroek et al., 1998; Dupree et al., 2005, 2020; Granzer et al., 2022).
+Large-scale convection is expected to be ubiquitous in RSG envelopes (e.g.
+Stothers, 1972; Goldberg et al., 2022; Chiavassa et al., 2011, 2024; Ma et al., 2025), will induce large-scale asymmetries in measurements of the stellar photocenter (Chiavassa et al., 2020) and spatially resolved RV measurements (Ma et al., 2024).
+If the convective activity seeds chromospheric activity, this should indeed impact Betelgeuse’s emission line profiles at the quadrant-by-quadrant and hemisphere-by-hemisphere level.
+Figure 7 also shows spectroscopic differences at the location of the H2 line at the companion’s expected RV, indicated by a vertical grey dotted line.
+While hints of a feature may appear in some quadrants, there is no clear evidence of emission lines from the companion at its expected RV offset within the measurement uncertainties in this feature.
+We now turn to discussing other stronger diagnostic lines.
 4.2 Emission from diagnostic lines
-Figure 7 highlights specific wavelength ranges containing diagnostic lines useful for inferring the presence of a companion (see Figure 2 and surrounding discussion in §2.3).
-These four wavelength ranges represent the most promising areas (other than Ly-, see §4.3) to detect the companion.
-We plot the NW (cyan) and SE (magenta) stacked hemispheres of our observations888We show halves instead of quadrants for ease in viewing, but the results are the same when considering all four quadrants, the Carpenter et al.
+While Ly- normally would be the most promising region to search for emission from a companion, the FUV spectrum shortward of 1235Å in Betelgeuse appears to show no flux above the noise level, potentially absorbed by Ly- damping wings in the outflowing wind.
+We note also that the feature near Ly- in Betelgeuse’s spectrum is primarily geocoronal (discussed in Appendix A).
+Given the varied nature of supergiant wind models (see van Loon 2025 for a recent review), it is difficult to assess the neutral absorption column along the line of sight to the putative companion’s orbital position.
+Thus we cannot determine whether the observed lack of emission in the Ly- region is due to absence of a YSO companion, or whether tangential absorption from Betelgeuse’s wind is significant enough to mask the presence of a companion with a relatively narrow Ly-alpha emission feature.
+Therefore, we turn to alternative wavelength regions containing promising diagnostic lines for inferring the presence of a companion (see Fig.
+2 and surrounding discussion in § 2.3).
+Examples are shown in Fig.
+8.
+We plot the NW (cyan) and SE (magenta) stacked hemispheres of our observations101010We show halves instead of quadrants for ease in viewing, but the results are the same when considering all four quadrants, the Carpenter et al.
 (2018) data (black), and several observations of YSOs from ULLYSES Roman-Duval et al.
@@
 These YSOs have been distance (168 pc Joyce et al.
-2020)999At the farther astrometric distance of 222pc (Harper et al., 2017), this would lead to a factor of reduction in flux.
-and extinction ( and MontargÚs et al.
-2021) corrected to the parameters of the Betelgeuse system.
-First, we compare the shapes of emission features between the NW and SE hemispheres, and to the previous 2011 Carpenter et al.
+2020)111111At the farther astrometric distance of 222pc (Harper et al., 2017), this would lead to a factor of reduction in flux.
+and extinction ( and , Montargès et al.
+2021) corrected to the parameters of the Betelgeuse system, as detailed in §2.3.
+We note that caution is needed when comparing the observed extinction applied here () to the discussion of the absorbing Hydrogen column () by O’Grady et al.
+(2025).
+Because the X-ray absorption column close to the stellar surface is unconstrained from direct X-ray observations, O’Grady et al.
+(2025) adopt estimates from stellar wind models of near the companion’s location.
+Even higher estimates of can be inferred from ALMA observations, e.g., integrating the profile determined in recent work by Dent et al.
+(2024).
+Applying the Predehl & Schmitt (1995) linear relationship between extinction and Hydrogen column for interstellar dust, , to even the low () or moderate () values of from O’Grady et al.
+(2025) yields implausibly high extinction values of or , respectively.
+In fact, the ISM relation does not apply to the partially-ionized circumstellar regime, where the dust-to-gas ratio is significantly smaller than it is in the farther-out ISM.
+A large hydrogen column in the star’s circumstellar wind environment could contribute to photoelectric absorption of X-rays, whereas a smaller amount of dust further out could modestly affect the FUV extinction.
+We compare the shapes of emission features between the NW and SE hemispheres, and to the previous 2011 Carpenter et al.
 (2018) Betelgeuse observations.
@@
 Given this, we look to the ULLYSES YSO data to illustrate what upper limits we can place on the mass and properties of the companion.
-We rule out the presence of a M2Mâ companion (light green line), which has flux greater than that of Betelgeuse in all cases.
-For the 0.77Mâ and 0.91Mâ YSOs (navy lines), even their chromospheric features with the highest flux relative to that of Betelgeuse (panel d) appear too low in flux to be discerned from noise in our spectra.
-For the 1.1Mâ and 1.5Mâ YSOs, the presence of an RV-shifted feature in just one of our observed quadrants/halves would be difficult to disentangle from the noise (see, e.g.
+We rule out the presence of a M2M⊙ companion (light green line), which has flux greater than that of Betelgeuse in all cases.
+For the 0.77M⊙ and 0.91M⊙ YSOs (navy lines), even their chromospheric features with the highest flux relative to that of Betelgeuse (panel d) appear too low in flux to be discerned from noise in our spectra.
+For the 1.1M⊙ and 1.5M⊙ YSOs, the presence of an RV-shifted feature in just one of our observed quadrants/halves would be difficult to disentangle from the noise (see, e.g.
 the CIV feature in panel d, as well as the CII doublet at 1334.5 and 1335.7 in panel a).
-In panels a), b), and c), for the narrow emission lines and not the disk emission features, we see that the flux of the 1.5Mâ YSO (e.g.
-the line near 1338.75 in panel a), or both that and the 1.1Mâ YSO (the CII lines near 1334.5-1336, H2 near 1505), meet or even slightly exceed that of all the Betelgeuse spectra.
-If is a 1.1-1.5Mâ YSO, and has a spectral profile similar to the example YSOs from ULLYSES, then we would rule out those masses, though the diversity in emission from YSOs renders these ULLYSES spectra imperfect templates for the spectrum of the companion.
+In panels a), b), and c), for the narrow emission lines and not the disk emission features, we see that the flux of the 1.5 YSO (e.g.
+the line near 1338.75Å in panel a), or both that and the 1.1 YSO (the CII lines near 1334.5-1336Å, H2 near 1505Å), meet or even slightly exceed that of all the Betelgeuse spectra.
+If is a 1.1-1.5M⊙ YSO, and has a spectral profile similar to the example YSOs from ULLYSES, then we would rule out those masses, though the diversity in emission from YSOs renders these ULLYSES spectra imperfect templates for the spectrum of the companion.
 Moreover, will be embedded within the dusty circumstellar material surrounding Betelgeuse, which may obscure or otherwise reprocess emission from the companion.
@@
 While our observational campaign was designed to be maximally sensitive to a possible detection, given the diversity inherent in YSO spectra in the FUV, a non-detection is unfortunately only weakly constraining.
-Take, for example, the strong difference in overall flux in emission features between the 0.91Mâ and 1.1Mâ YSO spectra.
+Take, for example, the strong difference in overall flux in emission features between the 0.91M⊙ and 1.1M⊙ YSO spectra.
 Differences of up to an order of magnitude are visible (e.g.
-panel a around 1336.5), despite only 0.2Mâ between the mass estimates.
-Therefore, we cannot conclusively rule out the presence of a 1Mâ companion, though a smaller (M1Mâ) is favored.
-4.3 Ly- Emission and Differences by quadrant
-The Ly- (1215.67) emission line is another strong feature in the UV chromosphere of Betelgeuse as well as other stars and YSOs (Linsky, 2017; Carpenter et al., 2018; Roman-Duval et al., 2020, 2025).
-However, there is significant contamination of this feature from geocoronal emission from Earthâs atmosphere.
-The only HST UV observations of the Ly- line in a star uncontaminated by geocoronal emission are from stars with extremely high (150 km/s) radial velocities (Guinan et al., 2016; Schneider et al., 2019).
-The 22 km/s radial velocity of the Betelgeuse system is within the approximate width (35 km/s) of the observed Ly- line.
-In an attempt to separate the source and geocoronal Ly- emission profiles, we extracted various cross-dispersion slices of the 2D spectra, including regions centered on the order traces where Ly- appears along with adjacent background regions on either side.
-Ly- appears in two adjacent orders in our E140M spectra, and for one of the orders the center trace spectrum shows evidence of a roughly 2.5 km s-1 shift relative to the background Ly- spectrum.
-However, no such signal was detected in the other order.
-In the left panel of Figure 8, we plot the four quadrants of our Betelgeuse observations (same colors as Figure 3) along with the 2011 ASTRAL observations in black, shifted to the RV of Betelgeuse at the time of our 2024 observations.
-We see no evidence of an additional, RV-shifted Ly- line in any of the four quadrants.
-The Ly- emission has the same general shape and width (0.1) in all quadrants.
-We do observe a small shift in the center wavelengths of the emission feature â the east quadrant is offset from the west quadrant by km/s, while the north and south are offset from the west by approximately half that value.
-The implied orientation â with the eastern side of Betelgeuse redshifted relative to the western side â is consistent with the rotation axis inferred by Dupree et al.
-(1987); Gilliland & Dupree (1996); Uitenbroek et al.
-(1998); Lobel & Dupree (2001); Kervella et al.
-(2009, 2018), and others.
-While large-scale convection ubiquitous in RSG envelopes (e.g.
-Stothers, 1972; Goldberg et al., 2022; Chiavassa et al., 2011, 2024; Ma et al., 2025) could over short timescales mimic both a changing astrometric signal (Chiavassa et al., 2020) and rotation inferred from an apparent redshift-blueshift dipole in resolved interferometric RV observations (Ma et al., 2024), convection is a stochastic process with decoherence times of order the envelope sound-crossing time (yr).
-The timing of our observations, on the other hand, has no correlation with the underlying convective phase or plume geometry of Betelgeuse.
-Therefore, the alignment of a south-westward RV dipole, persistent across many decades of observations and across many wavelengths including our 2024 observations, favors the claim that there is indeed coherent rotation of the near-surface layers.
-When comparing our observations to the 2011 observations of Carpenter et al.
-(2018), we also find a surprising km/s offset between our median Ly- measurements and the 2011 Ly- measurement.
-This is the only emission feature throughout the entire spectrum that displays this offset between the 2024 and 2011 observations, including the 1302-1306 OI lines, which are also contaminated by geocoronal emission.
-All other features are only offset by 2.5 km/s, the RV difference in Betelgeuse between these dates, not offset by this km/s blueshift.
-We cannot explain this offset, but note that our 2024 data do show Ly- at approximately the same RV as other features in the spectrum.
-Finally, in the right panel of Figure 8, we present our NW and SE spectra alongside several ULLYSES (Roman-Duval et al., 2020, 2025) YSOs, again distance and extinction corrected and shifted to âs RV at this phase.
-The YSOs plotted are different than in Figure 7; here we plot only ULLYSES YSOs observed with STIS, as COS is even more impacted by the geocoronal contamination at Ly- than STIS.
-Though we were not able to extract the true shape of the Ly- emission unique to Betelgeuse, we do note that the total RV of relative to Earth is 65 km/s (the vertical orange dotted line in Figure 8), and is beyond the extent of the Ly- line.
-As seen in the right panel of Fig.
-8, all reference YSO spectra (, , ) other than the 0.76Mâ YSO, the Ly- meet or exceed the flux of our Betelgeuse observations.
-While the lack of a feature at the location of the Ly- line for the RV of is clear, we cannot firmly conclude that Ly- emission from a YSO with mass Mâ would be visible above the noise level of our 2024 spectra, as the Ly- features of the ULLYSES data are themselves dominated by geocoronal contamination.
-If we nonetheless expect comparable intrinsic Ly- emission from YSOs, then the lack of a second emission line here is constraining.
-However, without uncontaminated emission profiles, it is impossible to place an exact mass constraint using Ly-.
+panel a around 1336.5Å), despite only 0.2M⊙ between the mass estimates.
+Therefore, we cannot conclusively rule out the presence of a 1M⊙ companion, though a smaller (M1M⊙) is favored by this analysis.
 4.3 Implications for the Companion
@@
 Figure 9 illustrates these scenarios with stellar evolutionary tracks (left panel) from MESA Isocrones and Stellar Tracks (MIST; Choi et al., 2016; Dotter, 2016; Paxton et al., 2011, 2013, 2015, 2018, left panel) at the masses given for the ULLYSES objects shown in Figure 8.
-The tracks shown span the Hayashi track through truncation at twice Betelgeuseâs approximate age, 20 Myr (prior to the zero-age main sequence in most cases).
+The tracks shown span the Hayashi track through truncation at twice Betelgeuse’s approximate age, 20 Myr (prior to the zero-age main sequence in most cases).
 Lines of constant age (isochrones, red) at 5 Myr and 15 Myr are shown, with thick curves between (purple-blue-green) liberally indicating approximate plausible ages of .
@@
 2005 and seen in our Figure 1), make it difficult to place strong constraints on the companion mass.
-Nevertheless, it is interesting that we only encounter true exclusion for a relatively high mass of (dotted green curve), and marginal constraints for the 1.1â1.5 cases (dashed blue-green curves).
-The fact that a YSO as massive as is not fully excluded leaves open the need for future resolved and pan-chromatic observing campaigns to place more detailed constraints on the companionâs physical nature and interaction with its environment, or, to identify a different cause of the Long Secondary Period persistent in Betelgeuseâs photometry, astrometry, and radial velocity from which the putative companion was inferred.
-Likewise, flux-calibrated libraries like XSL (Chen et al., 2014; Manara et al., 2017) extended to the FUV would be valuable for constraining quiet or near-MS companions in the range which arenât well-captured in the ULLYSES data set, as broader libraries across a wider parameter space and absolute flux calibration could facilitate a more realistic detection threshold.
-Either way, the most promising observational path forward for direct detection remains long-baseline observations and analysis of historical observations of the resolved stellar surface and circumstellar envelope, specifically informed by timing relative to the LSP phase.
-Likewise, future and proposed FUV missions such as the Large UV/Optical/IR Surveyor (LUVOIR, The LUVOIR Team 2019) and the Cosmological Advanced Survey Telescope for Optical and ultraviolet Research (CASTOR, CĂŽte et al.
+Nevertheless, it is interesting that we only encounter true exclusion for a relatively high mass of (dotted green curve), and marginal constraints for the 1.1–1.5 cases (dashed blue-green curves).
+The fact that a YSO as massive as is not fully excluded leaves open the need for future resolved and pan-chromatic observing campaigns to place more detailed constraints on the companion’s physical nature and interaction with its environment, or, to identify a different cause of the Long Secondary Period persistent in Betelgeuse’s photometry, astrometry, and radial velocity from which the putative companion was inferred.
+Likewise, flux-calibrated libraries like XSL (Chen et al., 2014; Manara et al., 2017) extended to the FUV would be valuable for constraining quiet or near-MS companions in the range which aren’t well-captured in the ULLYSES data set, as broader libraries across a wider parameter space and absolute flux calibration could facilitate a more realistic detection threshold.
+Either way, a promising observational path towards tighter constraints on the companion’s mass and surface proerties remains long-baseline observations and analysis of historical observations of the resolved stellar surface and circumstellar envelope, specifically informed by timing relative to the LSP phase.
+Likewise, future and proposed FUV missions such as the Large UV/Optical/IR Surveyor (LUVOIR, The LUVOIR Team 2019) and the Cosmological Advanced Survey Telescope for Optical and ultraviolet Research (CASTOR, Côte et al.
 2012) may provide additional, deeper constraints, in particular on the UV continuum emission of Betelgeuse as well as higher-contrast features from the companion at favorable orbital configurations.
+While this work was in the refereeing process, Howell et al.
+(2025) reported the probable claimed detection, via speckle imaging, of a source 6 magnitudes () fainter than Betelgeuse in the B band, with location and timing predicted by Goldberg et al.
+(2024) and MacLeod et al.
+(2025).121212Sadly, the measurement errors in HST (seen in Figure 7 and the right panel of Figure 9) are far too large to observe a differences in UV continuum flux across quadrants.
+The authors identify the companion as an young stellar object approaching but not yet on its Main Sequence.
+This scenario is prima facie at the margin of the range of YSO masses which we can exclude on the basis of a spectroscopic non-detection.
+However, chromospheric activity typically decreases as stars age closer to the Main Sequence (Preibisch & Feigelson, 2005; Linsky, 2017).
+Additionally, for a YSO companion interacting with Betelgeuse, spin-down processes may reduce the spin-controlled magnetic activity to below that of a typical pre-MS star.
+This may impact the emission limits on possible companion types.
+For example, since we expect a YSO companion to be stripped of any residual natal disk as discussed in §2.3, that may reduce any disk-locking that would otherwise spin up the companion prior to the disk-free T-Tauri stage.
+Similarly, tidal coupling between Betelgeuse and as discussed by MacLeod et al.
+(2025) would reduce the spin of the companion, as Betelgeuse’s spin period is slower than the companion’s orbital period.
+These scenarios both may suppress the companion’s spin and thus expected UV emission.
+Thus, the specific scenario discussed by Howell et al.
+(2025) is indeed cleanly within the constraints placed of our observations with HST.
+We therefore eagerly await upcoming efforts towards direct detection with speckle imaging, adaptive optics, and other high-resolution high-contrast observational techniques, in order to confirm the anticipated non-detection at occultation and detection at approaching quadrature.
 5 Conclusions
-In this work, we detailed our HST STIS E140M observations of Betelgeuse during mid-to-late November 2024, near its recently-predicted (Goldberg et al., 2024; MacLeod et al., 2025) companionâs receding limb passage.
-This resulted in a non-detection: we do not find any statistically significant line emission in the FUV which could be identified as originating from the companion at the expected RV offset from Betelgeuse (km/s redshift, corresponding to a AU orbit with a d period), within flux limits of ergâs-1âcm-2âĂ -1 which increase with increasing wavelength (across the full band, the average flux is ergâs-1âcm-2âĂ -1, and the error is ergâs-1âcm-2âĂ -1).
+In this work, we detailed our HST STIS E140M observations of Betelgeuse during mid-to-late November 2024, near its recently-predicted (Goldberg et al., 2024; MacLeod et al., 2025) companion’s receding limb passage.
+This resulted in a non-detection: we do not find any statistically significant line emission in the FUV which could be identified as originating from the companion at the expected RV offset from Betelgeuse (km/s redshift, corresponding to a AU orbit with a d period), within flux limits of erg s-1 cm-2 Å-1 which increase with increasing wavelength (across the full band, the average flux is erg s-1 cm-2 Å-1, and the error is erg s-1 cm-2 Å-1).
 There is likewise no statistically significant contribution to continuum emission from a companion (i.e.
-the continuum differences between quadrants are consistent with HSTâs error), and continuum-like emission from Betelgeuse is just barely at the level compared to the error.
+the continuum differences between quadrants are consistent with HST’s error), and continuum-like emission from Betelgeuse is just barely at the level compared to the error.
 Comparing to a set of YSO spectra taken with STIS in the ULLYSES database, we determine that higher-mass scenarios for the companion are ruled out, placing a spectroscopic mass constraint of and favoring masses , with acknowledgment of the great diversity in YSO spectra even within a given mass or spectral-type bin.
 This remains consistent with the lower-mass end of estimates fit from the STELLA RV only (Goldberg et al., 2024), but favors masses in the range as predicted by the long-baseline RV analysis (MacLeod et al., 2025).
-Although we cannot uniquely identify spectroscopic features from the companion itself, we see slight differences in the spectra for each quadrant of our HST observations which loosely correspond on the sky to the Northern, Southern, Eastern, and Western quadrants of Betelgeuseâs chromosphere and circumstellar envelope.
+Although we cannot uniquely identify spectroscopic features from the companion itself, we see slight differences in the spectra for each quadrant of our HST observations which loosely correspond on the sky to the Northern, Southern, Eastern, and Western quadrants of Betelgeuse’s chromosphere and circumstellar envelope.
 Though these differences are only slightly above the background, they likely indicate spatial variation across the chromosphere and circumstellar envelope.
-This is expected given the large-scale convective nature of RSG envelopes (see, e.g., recent review by Chiavassa et al., 2024), and is consistent with whatâs seen in interferometric studies of Betelgeuse (e.g.
-Harper et al., 2001; OâGorman et al., 2015; MontargĂšs et al., 2014, 2021; Kervella et al., 2009, 2018; Haubois et al., 2009, 2019, 2023; Kravchenko et al., 2021) as well as other work on spatially resolved UV observations of Betelgeuse and its chromosphere (e.g.
+This is expected given the scale of the vigorously convective nature of RSG envelopes (Stothers 1972, e.g.; Goldberg et al.
+2022, e.g.; Chiavassa et al.
+2009, 2011, 2024, e.g.; Ma et al.
+2025, e.g.).
+These asymmetries are also consistent with resolved interferometric observations of Betelgeuse (e.g.
+Harper et al., 2001; O’Gorman et al., 2015; Montargès et al., 2014, 2021; Kervella et al., 2009, 2018; Haubois et al., 2009, 2019, 2023; Kravchenko et al., 2021), as well as other work on spatially resolved UV observations of Betelgeuse and its chromosphere (e.g.
 Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Harper & Brown, 2006; Dupree et al., 2005, 2020; Matthews & Dupree, 2022).
-These differences are especially visible in the Ly- line (1215.67Ă ), though this feature is convolved with intense geocoronal emission and we were not able to cleanly separate the stellar Ly- profile.
-While the line itself is broadened by 10s of km/s, the velocity offsets of the Ly- line peaks in the four quadrants are consistent with the smeared-and-projected km/s rotational velocity (e.g.
-Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018).
-These spatial RV variations are also seen in the other emission features.
 Further modeling and future spatially-resolved observations may connect spectroscopic differences to spatial differences in chromospheric emission from Betelgeuse.
-Additional X-ray observations with Chandra, presented in OâGrady et al.
+Additional X-ray observations with Chandra, presented in O’Grady et al.
 (2025), exclude a compact object (white dwarf or neutron star), and constrain the X-ray emission coming from a possible YSO companion with larger uncertainties arising from uncertainty in the X-ray absorption along the line-of-sight as well as the multiple-orders-of-magnitude differences in X-ray emission from YSOs even of the same spectral class.
@@
 Acknowledgements
-We thank Morgan MacLeod for invaluable insight and discussions.
+We thank Morgan MacLeod for invaluable discussions.
 We likewise acknowledge helpful conversations with Andrea Antoni, Matteo Cantiello, Thavisha Dharmawardena, Zarina Dhillon, Kareem El-Badry, Brian Metzger, Michael Palumbo, Adiv Paradise, Mathieu Renzo, and Lieke van Son.
+We are forever grateful to Bill Paxton for all he has done for the astrophysical community.
 J.A.G.
@@
 M.J.
-gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commissionâs Widening Fellowship.
-This project has received funding from the European Unionâs Horizon 2020 research and innovation programme.
-This research was supported by the âSeismoLabâ KKP-137523 Ălvonal grant and by the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH).
+gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commission’s Widening Fellowship.
+This project has received funding from the European Union’s Horizon 2020 research and innovation programme.
+This research was supported by the ‘SeismoLab’ KKP-137523 Élvonal grant and by the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH).
 M.R.D.
@@
 This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium).
-This research made use of NASAâs Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.
+This research made use of NASA’s Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.
+Appendix A Nondetection of Lyman- Emission from companion
+The Ly- (1215.67Å) emission line is another often-shown feature in the FUV (Linsky, 2017; Carpenter et al., 2018; Roman-Duval et al., 2020, 2025).
+However, this feature is dominated by contamination from geocoronal emission from Earth’s atmosphere.We attempted, unsuccessfully, to separate possible source and geocoronal Ly- emission profiles by extracting cross-dispersion slices of the 2D spectra, including regions centered on the order traces where Ly- appears along with adjacent background regions on either side.
+Ly- appears in two adjacent orders in our E140M spectra.
+For one of the orders, the center trace spectrum shows evidence of a roughly 2.5 km s-1 shift relative to the background Ly- spectrum, but no such signal was detected in the other order.
+Nonetheless, the companion’s expected radial velocity shift is larger than the approximate width (35 km/s) of the geocoronal Ly- line, so if the companion were strongly emitting in Ly- it could in-principle be visible here.
+Therefore, for completeness, Figure 10 shows spectra near the Ly- feature from each of the four observed quadrants, along with the 2011 observations by Carpenter et al.
+(2018) shifted to the RV of Betelgeuse at the time of our 2024 observations.
+There is no discernable Ly- emission at the expected redshift of the companion.
+Due to the geocoronal origin of the strong Ly- features here, the km/s offset between the 2011 Ly- and the median 2024 Ly- can be attributed to differences in velocity about the Earth’s orbit projected toward Betelgeuse between the different epochs.
+The broad nature of the Ly- features is likewise attributed to the diffuse sky-glow filling the 0.2x0.2 aperture.
+Interestingly, Ly- is the only emission feature throughout the entire spectrum that displays this large of a velocity offset between the 2024 and 2011 observations; all other features are only offset by 2.5 km/s between the 2011 and our 2024 observations, owing to the RV difference in Betelgeuse between these dates.
+This includes the 1302-1306Å OI lines, which are also contaminated by geocoronal emission.
+We also compared with template spectra (not shown here) of YSO’s from ULYSSES (Roman-Duval et al., 2020, 2025), but those spectra are likewise contaminated by geocoronal emission.
+Without clean expectations for the uncontaminated Ly- emission profiles of possible companion scenarios, it is unfortunately impossible to place mass constraints from the lack of observed Ly- from the companion.